### v2

# Continuous-Time Quantum Markov Chains And Discretizations Of $p$-Adic Schrödinger Equations: Comparisons And Simulations

Abstract.
The continuous-time quantum walks (CTQWs) are a fundamental tool in the development of quantum algorithms. Recently, it was shown that discretizations of p-adic Schrödinger equations give rise to continuous-time quantum Markov chains (CTQMCs); this type of Markov chain includes the CTQWs constructed using adjacency matrices of graphs as a particular case. In this paper, we study a large class of p-adic Schrödinger equations and the associated CTQMCs by comparing them with p-adic heat equations and the associated continuous-time Markov chains (CTMCs). The comparison is done by a mathematical study of the mentioned equations, which requires, for instance, solving the initial value problems attached to the mentioned equations, and through numerical simulations. We conducted multiple simulations, including numerical approximations of the limiting distribution. Our simulations show that the limiting distribution of quantum Markov chains is greater than the stationary probability of their classical counterparts, for a large class of CTQMCs.
1. Introduction
Quantum computing is a computing paradigm that uses quantum physics effects like superposition, interference, and entanglement to achieve computational power beyond classical computing; see, e.g., the recent survey [1], and the references therein. The continuous-time quantum walks (CTQWs) are the quantum counterparts of the classical random walks. CTQWs are a fundamental tool for developing efficient quantum algorithms, simulating complex quantum systems, and potentially achieving quantum advantages in various computational tasks. CTQWs offer potential advantages over classical algorithms, particularly in graph-based problems like spatial search, [2]-[9]. Currently, there is a large variety of quantum physical systems that can be used for implementing quantum walks and their applications. For instance, trapped-ion and trapped-atom systems, bulk-optics systems, and integrated photonics systems. These considerable advances in physical implementations of quantum walks are allowing the implementation of specialized quantum walk computing systems for applications of practical interest, [1], [10].
Ultrametricity in physics means the emergence of ultrametric spaces in physical models. A metric space $(M,d)$ is ultrametric if the distance $d$ satisfies $d(A,B)\leq\max\left\{d(A,C),d(C,B)\right\}$ for any three points $A$, $B,$ $C$ in $M$. Ultrametricity was discovered in the 80s by Parisi and others in the theory of spin glasses and by Frauenfelder and others in physics of proteins; see e.g. [11], [12], see also [13, Chapter 4], and the references therein. In both cases, the space of states of a complex system has a hierarchical structure that plays a central role in the physical behavior of the system.
On the other hand, in the 1930s, Bronstein showed that general relativity and quantum mechanics imply that the uncertainty $\Delta x$ of any length measurement satisfies $\Delta x\geq L_{\text{Planck}}:=\sqrt{\frac{\hbar G}{c^{3}}}$, where $L_{\text{Planck}}\approx 10^{-33}$ cmcm is the Planck length. This implies that space-time is not an infinitely divisible continuum, [14]. Mathematically speaking, space-time must be a completely disconnected topological space. The ultrametric spaces are naturally completely disconnected. The field of $p$-adic numbers $\mathbb{Q}_{p}$ is a paramount example of an ultrametric space, with a very rich mathematical structure. There are several possible interpretations of the Bronstein inequality. One of them drives the loop quantum gravity. Volovich gave another interpretation of Bronstein’s inequality in the 80s. The inequality mentioned implies that real numbers cannot be used in models at the level of Planck’s length, because the Archimedean axiom, which appears naturally if we use real numbers, implies that lengths can be measured with arbitrary precision. Volovich proposed using $p$-adic numbers in physical models at the Planck scale, [15], see also [17], [16, Chapter 6].
In the Dirac-von Neumann formulation of quantum mechanics (QM), the states of a closed quantum system are vectors of an abstract complex Hilbert space $\mathcal{H}$, and the observables correspond to linear self-adjoint operators in $\mathcal{H}$. In $p$-adic QM, $\mathcal{H}=L^{2}(\mathbb{Q}_{p}^{N})$ or $\mathcal{H}=L^{2}(U)$, where $U\subset\mathbb{Q}_{p}^{N}$. This choice implies that the position vectors are $\mathbb{Q}_{p}^{N}$, while the time is a real variable. Using $\mathbb{R}\times\mathbb{Q}_{p}^{3}$ as a model of space-time, implies assuming that the space ($\mathbb{Q}_{p}^{3}$) has a discrete nature: there are no continuous word lines joining two different points in the space. The assumption of the discreteness of space, which requires passing from $\mathbb{R}\times\mathbb{R}^{3}$ to $\mathbb{R}\times\mathbb{Q}_{p}^{3}$ (or $\mathbb{R}\times\mathcal{X}$), is incompatible with special and general relativity. We warn the reader that there are several different types of $p$-adic QM; for instance, if the time is assumed to be a $p$-adic variable, the QM obtained radically differs from the one considered here. To the best of our knowledge, $p$-adic QM started in the 1980s under the influence of Vladimirov and Volovich, [18]. The literature about $p$-adic QM is pervasive, and here we cited just a few works, [18]-[26].
In [26], the first author showed that a large class of $p$-adic Schrödinger equations is the scaling limit of certain continuous-time quantum Markov chains (CTQMCs). Practically, a discretization of such an equation gives a CTQMC. As a practical result, we construct new types of continuous-time quantum walks (CTQWs) on graphs using two symmetric matrices. This construction includes, as a particular case, the CTQWs constructed using adjacency matrices. The final goal of the mentioned work was to contribute in the understanding of the role of the hypothesis of the discreteness of space in the foundations of quantum mechanics. The connection between $p$-adic QM and CTQWs show that $p$-adic QM has a physical meaning. $p$-Adic QM is a non-local theory because the Hamiltonians used are non-local operators, and consequently, spooky actions at a distance are allowed. But, this theory is not a mathematical toy. The paradigm asserting that the universe in not locally real, implies that $p$-adic QM allows realism.
In this work, we study a large class of $p$-adic Schrödinger equations and the associated CTQMCs by comparing them with $p$-adic heat equations and the associated CTMCs. This work is motivated by [5], where a particular class CTQWs was compared with their classical counterparts. Here, the comparison is done by a mathematical study of the mentioned equations, which requires, for instance, solving the initial value problems attached to the mentioned equations, and through numerical simulations. A $p$-adic heat equation is a continuous (or scaling) limit of a CTMC. This assertion has been verified for some classes of equations, see [26], which include the one considered here. But, for general $p$-adic heat equations, it is an open conjecture. By performing a Wick rotation on the heat equation, we obtain a Schrödinger equation, which is a continuous limit of a CTQMC. For general equations, this assertion is an open problem, but for the ones considered here, it was established in [26].
The equations mentioned can be well-approximated by systems of ordinary linear differential equations. With the matrices defining these systems, we construct random walks. We present a new approach to study CTQWs, which does not use adjacency matrices of graphs to construct random walks, but includes this construction as a particular case. Typically, the construction of CTQWs is based on graphs; we prefer using CTQMCs instead of CTQWs to emphasize that we do not use graphs.
We denote by $\mathbb{Z}_{p}$, where $p$ is a prime number, the unit ball of $\mathbb{Q}_{p}$, which can be visualize as an infinite rooted tree. By cutting this tree at level $l\geq 1$, we a finite rooted tree $G_{l}=\mathbb{Z}_{p}/p^{l}\mathbb{Z}_{p}$, with $l+1$ levels. The $p$-adic heat equations on the unit ball considered here have the form
$\left\{\begin{array}[c]{l}\frac{\partial}{\partial t}u\left(x,t\right)=J\left(\left|x\right|_{p}\right)\ast u\left(x,t\right)-u\left(x,t\right)={\displaystyle\int\limits_{\mathbb{Z}_{p}}}J\left(\left|x-y\right|_{p}\right)\left\{u\left(y,t\right)-u\left(x,t\right)\right\}dy\text{, }\\ \\ u\left(x,0\right)=u_{0}\left(x\right),\end{array}\right$.
where $x\in\mathbb{Z}_{p},t\geq 0$, and $J\left(\left|x\right|_{p}\right)\geq 0$ with $\int_{\mathbb{Z}_{p}}J\left(\left|y\right|_{p}\right)dy=1$. Here $dx$ is the Haar measure of $\mathbb{Q}_{p}$, and $\left|\cdot\right|_{p}$ is the $p$-adic norm. Here, for the sake of simplicity, we omit the definition of the function spaces The solution of the above initial problem is $u\left(x,t\right)=Z_{0}(x,t)\ast u_{0}\left(x\right)$, where $Z_{0}(x,t)$ is a heat kernel, which the transition density function for a Markov process in $\mathbb{Z}_{p}$, see Proposition 9.1 and Theorem 9.1, In Appendix B.
The above equation is a very particular example of a $p$-adic heat equation. In the general framework, the $p$-adic heat equations are deeply connected with models of complex hierarchical systems; they have been studied extensively in the last forty years; see, e.g. [13], [27]-[32], and the references therein.
Now, we construct a CTMC attached to the above $p$-adic heat equation. We decompose the unit ball as a disjoint union of ball of smaller radius:
$\mathbb{Z}_{p}={\displaystyle\bigsqcup\limits_{I\in G_{l}}}\left(I+p^{l}\mathbb{Z}_{p}\right).$
We denote by $\Omega\left(p^{l}\left|x-I\right|_{p}\right)$ the characteristic function of the ball $I+p^{l}\mathbb{Z}_{p}$. We now construct a CTMC, with states $I\in G_{l}$, by defining the transition probability between states as
$p_{r,v}(t)=p^{l}{\displaystyle\int\limits_{v+p^{l}\mathbb{Z}_{p}}}Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)dx.$
Furthermore, $\lim_{t\rightarrow\infty}p_{r,v}(t)=p^{-l}$, which means that the Markov chain admits a stationary probability distribution, cf. Theorem 10.1, in Appendix C.
By performing a Wick rotation in the heat equation, we obtain a Schrödinger equation:
$\left\{\begin{array}[c]{l}\mathrm{i}\frac{\partial}{\partial t}\Psi\left(x,t\right)=-J\left(\left|x\right|_{p}\right)\ast\Psi\left(x,t\right)+\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\geq 0\\ \\ \Psi\left(x,0\right)=\psi_{0}\left(x\right),\end{array}\right$.
where $\mathrm{i}=\sqrt{-1}$. The solution of is given by $\Psi\left(x,t\right)=Z_{0}\left(x,it\right)\ast\psi_{0}\left(x\right)$, cf. Theorem 11.1, in Appendix C. Now, we construct a quantum Markov chain in $G_{l}$ using the following transition probabilities:
$\pi_{r,v}\left(t\right)=\left|\left\langle p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right),Z_{0}(x,it)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\right\rangle\right|^{2},$
cf. Proposition 11.1 and Theorem 11.2, in Appendix D.
Set $\boldsymbol{J}\varphi=J\ast\varphi-\varphi$, then the operators $\boldsymbol{J}$, $\mathrm{e}^{t\boldsymbol{J}}$, $\mathrm{e}^{it\boldsymbol{J}}$ are limits of finite dimensional operators $\boldsymbol{J}^{\left(l\right)}$, $\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}$, $\mathrm{e}^{it\boldsymbol{J}^{\left(l\right)}}$, and consequently in the computation of $p_{r,v}(t)$, respectively of $\pi_{r,v}\left(t\right)$, we can use $Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)=\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)$, respectively $Z_{0}(x,it)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)=\mathrm{e}^{it\boldsymbol{J}^{\left(l\right)}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)$, and thus the construction of the above-mentioned Markov chains depends on a matrix $\boldsymbol{J}^{\left(l\right)}$. By suitably selecting this matrix, we construct the CTQWs based on adjacency matrices of graphs, see Appendix E.
As in [5], [3], the comparison between CTQMCs and CTMCs is based on the comparison between
$p_{I}^{\text{sta}}=\lim_{t\rightarrow\infty}p_{I,J}\left(t\right)=p^{-l},$
and the limiting distribution (or longtime average), which is defined as
$\chi_{I,J}=\lim_{T\rightarrow\infty}\frac{1}{T}{\displaystyle\int\limits_{0}^{T}}\pi_{I,J}\left(t\right)dt.$
We obtained a formula for $\pi_{I,J}\left(t\right)$ as a $p$-adic oscillatory integral, see Theorem 11.2, in Appendix D. But, this result is not sufficient to compute an explicit formula for $\chi_{I,J}$. We select a family of kernels $J_{\alpha}\left(\left|x\right|_{p}\right),\alpha\in\left(0,\infty\right),$ (the Bessel potentials), and study the properties of the CTQMCs and CTMCs via numerical simulations. Our simulations show that $\chi_{I,J}>p_{I}^{\text{sta}}$. This fact is interpreted as the computational power of the CTQMC is greater than that of the corresponding CTMC. All the matrices in our simulations have an ultrametric structure, which is natural since we are using $p$-adic numbers. The simulations given in [3] show that the matrices $\left[\chi_{I,J}\right]$ have an ultrametric structure supporting the generality of our approach.
Currently, we cannot give specific insights on how our techniques can be used in the development of quantum algorithms. However, as we mentioned in [26], the connection between $p$-adic QM and CTQMCs opens several new research problems. Here, we pose the question: what type of quantum networks are produced by discretizing non-linear $p$-adic Schrödinger equations? In our opinion, this question is deeply connected with the construction of quantum neural networks, see [33]-[34].
This work is organized into two parts. The first discusses the results and the simulations using a minimal mathematical background. The technical results were placed in appendices at the end of the paper.
2. Preliminaries
2.1. The field of $p$-adic numbers
From now on, we use $p$ to denote a fixed prime number. Any non-zero $p$-adic number $x$ has a unique expansion of the form
(2.1) $x=x_{-k}p^{-k}+x_{-k+1}p^{-k+1}+\ldots+x_{0}+x_{1}p+\ldots,\text{ }$
with $x_{-k}\neq 0$, where $k$ is an integer, and the $x_{j}$s are numbers from the set $\left\{0,1,\ldots,p-1\right\}$. The set of all possible sequences of the form (2.1) constitutes the field of $p$-adic numbers $\mathbb{Q}_{p}$. There are natural field operations, sum and multiplication, on series of form (2.1). There is also a norm in $\mathbb{Q}_{p}$ defined as $\left|x\right|_{p}=p^{-ord(x)}$, where $ord_{p}(x)=ord(x)=k$, for a nonzero $p$-adic number $x$. By definition $ord(0)=\infty$. The field of $p$-adic numbers with the distance induced by $\left|\cdot\right|_{p}$ is a complete ultrametric space. The ultrametric property refers to the fact that $\left|x-y\right|_{p}\leq\max\left\{\left|x-z\right|_{p},\left|z-y\right|_{p}\right\}$ for any $x$, $y$, $z$ in $\mathbb{Q}_{p}$. The $p$-adic integers which are sequences of form (2.1) with $k\geq 0$. All these sequences constitute the unit ball $\mathbb{Z}_{p}$. The unit ball is an infinite rooted tree with fractal structure. As a topological space $\mathbb{Q}_{p}$ is homeomorphic to a Cantor-like subset of the real line, see, e.g., [19], [35].
A function $\varphi:\mathbb{Q}_{p}\rightarrow\mathbb{C}$ is called locally constant, if for any $a\in\mathbb{Q}_{p}$, there is an integer $l=l(a)$, such that
$\varphi\left(a+x\right)=\varphi\left(a\right)\text{ for any }|x|_{p}\leq p^{l}.$
The set of functions for which $l=l\left(\varphi\right)$ depends only on $\varphi$ form a $\mathbb{C}$-vector space denoted as $\mathcal{U}_{loc}\left(\mathbb{Q}_{p}\right)$. We call $l\left(\varphi\right)$ the exponent of local constancy. If $\varphi\in\mathcal{U}_{loc}\left(\mathbb{Q}_{p}\right)$ has compact support, we say that $\varphi$ is a test function. We denote by $\mathcal{D}(\mathbb{Q}_{p})$ the complex vector space of test functions. There is a natural integration theory so that $\int_{\mathbb{Q}_{p}}\varphi\left(x\right)dx$ gives a well-defined complex number. The measure $dx$ is the Haar measure of $\mathbb{Q}_{p}$. In Appendix A, we give a quick review of the basic aspects of the $p$-adic analysis required here.
If $U$ is an open subset of $\mathbb{Q}_{p}$, $\mathcal{D}(U)$ denotes the $\mathbb{C}$-vector space of test functions with supports contained in $U$, then $\mathcal{D}(U)$ is dense in
$L^{\rho}\left(U\right)=\left\{\varphi:U\rightarrow\mathbb{C};\left\|\varphi\right\|_{\rho}=\left\{{\displaystyle\int\limits_{U}}\left|\varphi\left(x\right)\right|^{\rho}dx\right\}^{\frac{1}{\rho}}<\infty\right\},$
for $1\leq\rho<\infty$, see, e.g., [35, Section 4.3]. In this paper, we consider QM in the sense of the Dirac-von Neumann formulation on the Hilbert space $L^{2}\left(U\right)$. Given $f,g\in L^{2}\left(U\right)$, we set
$\left\langle f,g\right\rangle={\displaystyle\int\limits_{U}}f\left(x\right)\overline{g\left(x\right)}dx,$
where the bar denotes the complex conjugate.
Remark 2.1.
The above formulation and results can be easily extended to the $N$-dimensional space $\mathbb{Q}_{p}^{N}$. We extend the $p$-adic norm to $\mathbb{Q}_{p}^{N}$ by taking
$x||_{p}:=\max_{1\leq i\leq N}|x_{i}|_{p},\qquad\text{for }x=(x_{1},\dots,x_{N})\in\mathbb{Q}_{p}^{N}.$
If we define $ord(x)=\min_{1\leq i\leq N}\{ord(x_{i})\}$, then $x||_{p}=p^{-ord(x)}$. The metric space $\left(\mathbb{Q}_{p}^{N},||\cdot||_{p}\right)$ is a complete ultrametric space. Ultrametric means that the norm $||\cdot||_{p}$ satisfies
[…]
A function $\varphi:\mathbb{Q}_{p}^{N}\rightarrow\mathbb{C}$ is called locally constant, if for any $a\in\mathbb{Q}_{p}^{N}$, there is an integer $l=l(a)$, such that $\varphi\left(a+x\right)=\varphi\left(a\right)$ for any $x||_{p}\leq p^{l}$. We denote by $\mathcal{D}(\mathbb{Q}_{p}^{N})$ the complex vector space of test functions. There is a natural integration theory so that $\int_{\mathbb{Q}_{p}^{N}}\varphi\left(x\right)d^{N}x$ gives a well-defined complex number. The measure $d^{N}x$ is the Haar measure of $\mathbb{Q}_{p}^{N}$.
2.2. Additive characters
Using that any non-zero $p$-adic number $x$ admits an expansion of the form $x=p^{ord(x)}\sum_{j=0}^{\infty}x_{j}p^{j}$, $x_{0}\neq 0$, the fractional part of $x\in\mathbb{Q}_{p}$, denoted $\{x\}_{p}$, is the rational number defined as
$\left\{x\right\}_{p}=\left\{\begin{array}[c]{lll}0&\text{if}&x=0\text{ or }ord(x)\geq 0\\ &&\\ p^{ord(x)}\sum_{j=0}^{-ord(x)-1}x_{j}p^{j}&\text{if}&ord(x)<0.\end{array}\right$.
Set $\chi_{p}(y)=\exp(2\pi i\{y\}_{p})$ for $y\in\mathbb{Q}_{p}$. The map $\chi_{p}(\cdot)$ is an additive character on $\mathbb{Q}_{p}$, i.e., a continuous map from $\left(\mathbb{Q}_{p},+\right)$ into $S$ (the unit circle considered as a multiplicative group) satisfying $\chi_{p}(x_{0}+x_{1})=\chi_{p}(x_{0})\chi_{p}(x_{1})$, $x_{0},x_{1}\in\mathbb{Q}_{p}$. The additive characters of $\mathbb{Q}_{p}$ form an Abelian group which is isomorphic to $\left(\mathbb{Q}_{p},+\right)$. The isomorphism is given by $\xi\rightarrow\chi_{p}(\xi x)$, see, e.g., [35, Section 2.3].
2.3. $p$-Adic QM
In the Dirac-Von Neumann formulation of QM, to every isolated quantum system there is associated a separable complex Hilbert space $\mathcal{H}$ called the space of states. The Hilbert space of a composite system is the Hilbert space tensor product of the state spaces associated with the component systems. Non-zero vectors describe the states of a quantum system from $\mathcal{H}$. Each observable corresponds to a unique linear self-adjoint operator in $\mathcal{H}$. The most important observable of a quantum system is its energy. We denote the corresponding operator by $\boldsymbol{H}$. Let $\Psi_{0}\in\mathcal{H}$ be the state at time $t=0$ of a certain quantum system. Then at time $t$ the system is represented by the vector $\Psi\left(t\right)=\boldsymbol{U}_{t}\Psi_{0}$, where $\boldsymbol{U}_{t}=\mathrm{e}^{-it\boldsymbol{H}}$, $t\geq 0$, is a unitary operator called the evolution operator. The vector function $\Psi\left(t\right)$ is differentiable if $\Psi\left(t\right)$ is contained in the domain $Dom(\boldsymbol{H})$ of $\boldsymbol{H}$, which happens if at $t=0$, $\Psi_{0}\in Dom(\boldsymbol{H})$, and in this case the time evolution of $\Psi\left(t\right)$ is controlled by the Schrödinger equation $\mathrm{i}\frac{\partial}{\partial t}\Psi\left(t\right)=\boldsymbol{H}\Psi\left(t\right)$, where $\mathrm{i}=\sqrt{-1}$ and the Planck constant is assumed to be one. For an in-depth discussion of QM, the reader may consult [36]-[37]. In $p$-adic QM, $\mathcal{H=}L^{2}\left(U\right)$, where $U$ is an open compact subset of $\mathbb{Q}_{p}$. For the sake of simplicity, we use $\mathcal{H=}L^{2}\left(\mathbb{Z}_{p}\right)$, where $\mathbb{Z}_{p}=\left\{x\in\mathbb{Q}_{p};\left|x\right|_{p}\leq 1\right\}$ is the unit ball of $\mathbb{Q}_{p}$; $\mathbb{Z}_{p}$ is a compact additive group.
3. CTQMCs and $p$-adic Schrödinger equations
In this section, we discuss two distinct techniques for constructing CTQMCs based on $p$-adic Schrödinger equations. Most constructions of CTQWs are based on the use of adjacency matrices of graphs. This approach cannot be used to construct the CTQMCs introduced in this section. On the contrary, most CTQWs based on graphs can be constructed as a discretization of suitable $p$-adic Schrödinger equations; see [26], and Appendix E.
3.1. Initial value problems for $p$-adic Schrödinger equations on the unit ball
We set $\boldsymbol{H}:\mathcal{D}(\mathbb{Z}_{p})\rightarrow L^{2}(\mathbb{Z}_{p})$ to be a self-adjoint operator defined on the dense subspace $\mathcal{D}(\mathbb{Z}_{p})\hookrightarrow L^{2}(\mathbb{Z}_{p})$. Consider the following Cauchy problem
(3.1) $\left\{\begin{array}[c]{l}\Psi\left(\cdot,t\right)\in L^{2}(\mathbb{Z}_{p})\text{, }t>0;\text{ }\Psi\left(x,\cdot\right)\in C^{1}(0,\infty)\text{, }x\in\mathbb{Z}_{p};\\ \\ \mathrm{i}\frac{\partial}{\partial t}\Psi\left(x,t\right)=\boldsymbol{H}\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t>0;\\ \\ \Psi\left(x,0\right)=\Psi_{0}\left(x\right)\in L^{2}(\mathbb{Z}_{p})\text{;}\\ \\ \left\|\Psi\left(\cdot,t\right)\right\|_{2}=\left\|\Psi_{0}\right\|_{2}=1\text{, }t\geq 0.\end{array}\right$.
We assume the existence of a unique solution for (3.1). We denote the unitary group attached to $\boldsymbol{H}$ as $\mathrm{e}^{-i\boldsymbol{H}t}$, $t\in\mathbb{R}$.
3.2. An orthonormal basis for $L^{2}(\mathbb{Z}_{p})$
We set
$\mathbb{Q}_{p}/\mathbb{Z}_{p}=\left\{\sum_{j=-1}^{-m}x_{j}p^{j};\text{for some }m>0\right\}.$
For $b\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$, $r\in\mathbb{Z}$, we denote by $\Omega\left(\left|p^{r}x-b\right|_{p}\right)$ the characteristic function of the ball $bp^{-r}+p^{-r}\mathbb{Z}_{p}$.
We now define
(3.2) $\Psi_{rbk}\left(x\right)=p^{\frac{-r}{2}}\chi_{p}(p^{-1}k\left(p^{r}x-b\right))\Omega\left(\left|p^{r}x-b\right|_{p}\right),$
where $r\in\mathbb{Z}$, $k\in\{1,\dots,p-1\}$, and $b\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$. Then,
(3.3) ${\displaystyle\int\limits_{\mathbb{Q}_{p}}}\Psi_{rbk}\left(x\right)dx=0,$
and $\left\{\Psi_{rbk}\left(x\right)\right\}_{rbk}$ forms a complete orthonormal basis of $L^{2}(\mathbb{Q}_{p})$; see, e.g., [35, Theorems 9.4.5 and 8.9.3], or [13, Theorem 3.3].
The set of functions
${\displaystyle\bigcup}\left\{\Omega\left(\left|x\right|_{p}\right)\right\}\text{ \ \ }{\displaystyle\bigcup}\text{ \ }{\displaystyle\bigcup\limits_{k\in\{1,\dots,p-1\}}}\text{ \ \ }{\displaystyle\bigcup\limits_{r\leq 0}}\text{ \ \ }{\displaystyle\bigcup\limits_{\begin{subarray}{c}bp^{-r}\in\mathbb{Z}_{p}\\ b\in\mathbb{Q}_{p}/\mathbb{Z}_{p}\end{subarray}}}\left\{\Psi_{rbk}\left(x\right)\right\},$
is an orthonormal basis in $L^{2}(\mathbb{Z}_{p})$, [38, Proposition 2].
We assume that $\Omega\left(\left|x\right|_{p}\right)$, and the $\Psi_{rbk}\left(x\right)$ are eigenfunctions of $\boldsymbol{H}$:
$\boldsymbol{H}\Omega\left(\left|x\right|_{p}\right)=E_{\text{gnd}}\Omega\left(\left|x\right|_{p}\right)\text{, }\boldsymbol{H}\Psi_{rbk}\left(x\right)=E_{rbk}\Psi_{rbk}\left(x\right)\text{ for }r,b,k,$
where $E_{\text{gnd}}$, $E_{rbk}$ are the corresponding eigenvalues. Then,
(3.4) $\Psi\left(x,t\right)=A_{0}\Omega\left(\left|x\right|_{p}\right)\mathrm{e}^{-iE_{\text{gnd}}t}+{\displaystyle\sum\nolimits_{r,b,k}}A_{rbk}\Psi_{rbk}\left(x\right)\mathrm{e}^{-iE_{rbk}t},$
where the complex constants $A_{0}$, $A_{rbk}$ are determined by the condition
(3.5) $\Psi\left(x,0\right)=\Psi_{0}\left(x\right)=A_{0}\Omega\left(\left|x\right|_{p}\right)+{\displaystyle\sum\nolimits_{r,b,k}}A_{rbk}\Psi_{rbk}\left(x\right).$
3.3. Construction of CTQMCs I
Let $\mathcal{N}\subset\mathbb{Z}_{p}$ be a subset of measure zero, and let $\mathcal{K}_{j}$ $\subset\mathbb{Z}_{p}$ be disjoint, open, compact subsets such that
(3.6) $\mathbb{Z}_{p}\smallsetminus\mathcal{N}={\displaystyle\bigsqcup\limits_{j\in\mathbb{J}}}\mathcal{K}_{j},$
where $\mathbb{J}$ is countable (i.e., $\mathbb{J}$ is either finite or countably infinite).
We set $e_{v}=c_{v}1_{\mathcal{K}_{v}}$, where $1_{\mathcal{K}_{v}}$ is the characteristic function of $\mathcal{K}_{v}$, and the constant $c_{v}$ is chosen so that $\left\|e_{v}\right\|_{2}=1$. Then, $\left\{e_{j}\right\}_{j\in\mathbb{J}}$ is an orthonormal subset of $L^{2}(\mathbb{Z}_{p})$. We assume that
(3.7) $\mathcal{H}_{\mathbb{J}}:=\mathit{Span}\left\{e_{j};j\in\mathbb{J}\right\}\text{ is a Hilbert subspace of }L^{2}(\mathbb{Z}_{p}),$
and that
(3.8) $\mathrm{e}^{-\mathrm{i}\boldsymbol{H}t}\mathcal{H}_{\mathbb{J}}\subset\mathcal{H}_{\mathbb{J}}\text{.}$
The function
(3.9) $\Psi_{v}\left(x,t\right):=\mathrm{e}^{-\mathrm{i}\boldsymbol{H}t}e_{v}(x)\in L^{2}\left(\mathbb{Z}_{p}\right),$
is the solution of the Cauchy problem (3.1), with $\Psi_{0}\left(x\right)=e_{v}(x)$.
We construct a network with nodes $j\in\mathbb{J}$, where $j$ is identified with $1_{\mathcal{K}_{j}}$, and define the transition probability between nodes $v$ and $r$ as
(3.10) $\pi_{r,v}\left(t\right)=\left|\left\langle e_{r},\mathrm{e}^{-\mathrm{i}\boldsymbol{H}t}e_{v}\right\rangle\right|^{2}=\left|\left\langle e_{r},\Psi_{v}\right\rangle\right|^{2},$
where $\left\langle\cdot,\cdot\right\rangle$ is the inner product of $L^{2}\left(\mathbb{Z}_{p}\right)$. Now, the Cauchy-Schwarz inequality,
$0\leq\sqrt{\pi_{r,v}\left(t\right)}\leq\left\|e_{r}\right\|_{2}\left\|\Psi_{v}\right\|_{2}=\left\|e_{r}\right\|_{2}\left\|e_{v}\right\|_{2}=1,$
implies that $\pi_{r,v}\left(t\right)\in\left[0,1\right]$.
Proposition 3.1.
Under the hypotheses (3.7) and (3.8), ${\textstyle\sum\nolimits_{r\in\mathbb{J}}}\pi_{r,v}\left(t\right)=1$, for $t\geq 0$, and any $v\in\mathbb{J}$.
Proof.
First,
$L^{2}(\mathbb{Z}_{p})=\mathcal{H}_{\mathbb{J}}{\textstyle\bigoplus}\mathcal{H}_{\mathbb{J}}^{\perp}\text{,}$
where $\mathcal{H}_{\mathbb{J}}^{\perp}$ is the orthogonal complement of $\mathcal{H}_{\mathbb{J}}$. Then, for each $t\geq 0$,
$\Psi_{v}\left(x,t\right)={\displaystyle\sum\limits_{r\in\mathbb{J}}}c_{r}\left(t\right)e_{r}\left(x\right)+z_{v}(x,t),$
where $z_{v}(x,t)\in\mathcal{H}_{\mathbb{J}}^{\perp}$, and
(3.11) $\left\|\Psi_{v}\left(\cdot,t\right)\right\|_{2}^{2}={\displaystyle\sum\limits_{r\in\mathbb{J}}}\left|c_{r}\left(t\right)\right|^{2}+\left\|z_{v}(\cdot,t)\right\|_{2}^{2}=1.$
On the other hand, by using that $\pi_{r,v}\left(t\right)=\left|\left\langle e_{r}\left(x\right),\Psi_{v}\left(x,t\right)\right\rangle\right|^{2}=\left|c_{r}\left(t\right)\right|^{2}$, and (3.11),
${\textstyle\sum\nolimits_{r\in\mathbb{J}}}\pi_{r,v}\left(t\right)={\textstyle\sum\nolimits_{r\in\mathbb{J}}}\left|c_{r}\left(t\right)\right|^{2}=1\Leftrightarrow\text{ }z_{v}(x,t)=0\text{ }\Leftrightarrow\Psi_{v}\left(x,t\right)\in\mathcal{H}_{\mathbb{J}},$
for any $t\geq 0$. ∎
In the framework of continuous-time Markov chains, $\pi_{s,r}\left(t\right)$ represents the probability that the chain is in state $s$ at time $t$, given it started in state $r$ at time 0.
Theorem 3.1.
Under the hypotheses 3.7 and 3.8, $\left[\pi_{s,r}\left(t\right)\right]_{s,r\in\mathbb{J}}$ is the transition matrix of a continuous-time quantum Markov chain with state space $\mathbb{J}$.
We now study the case when hypothesis (3.8) is false.
Proposition 3.2.
Assume that hypothesis (3.7) is true, and that hypothesis (3.8) is false. Then, ${\textstyle\sum\nolimits_{r\in\mathbb{J}}}\pi_{r,v}\left(t\right)=1$ if and only if $\left\{e_{j}\right\}_{j\in\mathbb{J}}$ is a complete orthonormal subset (a basis) of $L^{2}(\mathbb{Z}_{p})$. Furthermore, if $\mathbb{J}$ is a finite set, ${\textstyle\sum\nolimits_{r\in\mathbb{J}}}\pi_{r,v}\left(t\right)<1$.
Proof.
By reasoning as in the proof of the previous proposition, we have
${\textstyle\sum\nolimits_{r\in\mathbb{J}}}\pi_{r,v}\left(t\right)={\textstyle\sum\nolimits_{r\in\mathbb{J}}}\left|c_{r}\left(t\right)\right|^{2}=1\text{ if and only if }z_{v}(x,t)=0,$
for any $t\geq 0$. The fact $z_{v}(x,t)=0$ is equivalent to the set $\left\{e_{j}\right\}_{j\in\mathbb{J}}$ is complete in $L^{2}(\mathbb{Z}_{p})$. The last assertion follows from the fact that $L^{2}(\mathbb{Z}_{p})$ is an infinite-dimensional Hilbert space. ∎
With the hypotheses of the above proposition, we assume that $\mathbb{J}$ is a finite set. We attach to each $e_{r}$, $r\in\mathbb{J}$, a state, denote as $r$, and define a transition probability $\pi_{r,v}\left(t\right)$ for $r$, $v\in\mathbb{J}$, as in (3.10). We introduce an extra state denoted as $\infty$, and define
$\pi_{\infty,v}\left(t\right)=1-{\textstyle\sum\nolimits_{r\in\mathbb{J}}}\pi_{r,v}\left(t\right)\text{, for }v\in\mathbb{J},$
and
$\pi_{s,\infty}\left(t\right)=\left\{\begin{array}[c]{ccc}0&\text{if}&s\in\mathbb{J}\\ &&\\ 1&\text{if}&s=\infty,\end{array}\right$.
for any $t\geq 0$.
Theorem 3.2.
Assume that hypothesis (3.7) is true, and that $\mathbb{J}$ is a finite set (which means that the hypothesis (3.8) is false). Then, $\left[\pi_{s,r}\left(t\right)\right]_{s,r\in\mathbb{J}\cup\left\{\infty\right\}}$ is the transition matrix of a continuous-time quantum Markov chain with state space $\mathbb{J}\cup\left\{\infty\right\}$.
The presented construction is entirely different from the one given in [2]-[5], since here we use an infinite-dimensional Hilbert space $L^{2}(\mathbb{Z}_{p})$, and we do not use graphs. In contrast, in the references mentioned above, the constructions use finite-dimensional Hilbert spaces and adjacency matrices. On the other hand, we propose to interpret the state $\infty$ in the previous theorem as an interaction with an external system. We do not present numerical simulations for this type of network here. The main difficulty is the computational implementation of the condition $\mathrm{e}^{-\mathrm{i}\boldsymbol{H}t}\mathcal{H}_{\mathbb{J}}\nsubseteqq\mathcal{H}_{\mathbb{J}}$.
3.4. Construction of CTQMCs II
The second construction is based on the Born interpretation of the wavefunctions, which is also valid in the $p$-adic framework, [24]-[25]. Let $\mathcal{K}$ be an open compact subset of $\mathbb{Z}_{p}$ as before. The function $\left|\Psi_{v}\left(x,t\right)\right|^{2}$ is the probability density for the position of a particle/walker at the time $t$ given that at $t=0$ the particle/walker was in $\mathcal{K}_{v}$. By the Born interpretation
$\widetilde{\pi}_{r,v}\left(t\right):={\displaystyle\int\limits_{\mathcal{K}_{r}}}\left|\Psi_{v}\left(x,t\right)\right|^{2}dx$
is the probability of finding the particle/walker in $\mathcal{K}_{r}$; which we interpret as a transition probability from a $\mathcal{K}_{v}$ to $\mathcal{K}_{r}$. Now, using partition (3.6), we have
$1$ $={\displaystyle\int\limits_{\mathcal{K}}}\left|\Psi_{v}\left(x,t\right)\right|^{2}dx={\displaystyle\int\limits_{\mathcal{K}\smallsetminus\mathcal{N}}}\left|\Psi_{v}\left(x,t\right)\right|^{2}dx$
$={\displaystyle\sum\limits_{j\in\mathbb{J}}}\text{ }{\displaystyle\int\limits_{\mathcal{K}_{j}}}\left|\Psi_{v}\left(x,t\right)\right|^{2}dx={\displaystyle\sum\limits_{j\in\mathbb{J}}}\widetilde{\pi}_{j,v}\left(t\right).$
So, the matrix $\left[\widetilde{\pi}_{r,v}\left(t\right)\right]_{r,v\in\mathbb{J}}$ defines a quantum Markov chain with space state $\mathbb{J}$, where a walker jumps between the sets $\mathcal{K}_{v}$ to $\mathcal{K}_{r}$ with a transition probability $\widetilde{\pi}_{r,v}\left(t\right)$.
The computation of the transition probabilities $\widetilde{\pi}_{r,v}\left(t\right)$ is involved, even in dimension one and for simple partitions of the unit ball, [39].
4. $p$-Adic heat equations and CTMCs
In this section, we provide a brief discussion of the results in Appendix B, assuming a minimal mathematical background without including proofs. In the mentioned appendix, a rigorous treatment of the results of this section is given.
4.1. A class of $p$-adic heat equations
In [40], and the references therein, the first author established that CTMCs can be obtained as discretizations of p-adic heat equations. Here, based on this reference, we study a class of $p$-adic heat equations and the associated CTMCs. The techniques used here are different from the ones used in [40]. We fix a real-valued radial function $J(x)=J(\left|x\right|_{p})$, where $J:\left[0,1\right]\rightarrow\left[0,\infty\right)$, satisfying
$\left\|J(\left|x\right|_{p})\right\|_{1}={\displaystyle\int\limits_{\mathbb{Z}_{p}}}J(\left|x\right|_{p})dx=1.$
We extend $J(\left|x\right|_{p})$ as zero outside of the unit ball, then $J(\left|x\right|_{p})\in L^{1}\left(\mathbb{Q}_{p}\right)$, and consequently, the Fourier transform $\widehat{J}(\xi):\mathbb{Q}_{p}\rightarrow\mathbb{C}$ is well-defined. The Fourier transform does not preserve the support of $J(\left|x\right|_{p})$, i.e., $\widehat{J}(\xi)\neq 0$ for points $\xi$ outside of the unit ball.
We now construct an operator
$\begin{array}[c]{cccc}\boldsymbol{J}:&L^{\rho}\left(\mathbb{Z}_{p}\right)&\rightarrow&L^{\rho}\left(\mathbb{Z}_{p}\right)\\ &&&\\ &\varphi\left(x\right)&\rightarrow&J\left(\left|x\right|_{p}\right)\ast\varphi\left(x\right)-\varphi\left(x\right)\end{array}$
which is a linear, bounded operator, for $\rho\in\left[0,\infty\right]$. Here, we require only the cases $\rho=1,2$. Furthermore, $\boldsymbol{J}$ is a pseudo-differential operator,
$\boldsymbol{J}\varphi\left(x\right)=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\left(\widehat{J}(\left|\xi\right|_{p})-1\right)\mathcal{F}_{x\rightarrow\xi}\varphi\right)={\displaystyle\int\limits_{\mathbb{Q}_{p}}}\left(\widehat{J}(\left|\xi\right|_{p})-1\right)\widehat{\varphi}\left(\xi\right)\chi_{p}\left(-\xi x\right)dx,$
for $\varphi\in\mathcal{D}\left(\mathbb{Z}_{p}\right)$, where $\mathcal{F}$ denotes the Fourier transform, see Appendix A.
We fix a time horizon $T\in\left(0,\infty\right]$, and consider the following initial value problem:
(4.1) $\left\{\begin{array}[c]{l}\frac{\partial}{\partial t}u\left(x,t\right)=J\left(\left|x\right|_{p}\right)\ast u\left(x,t\right)-u\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\in\left[0,T\right]\\ \\ u\left(x,0\right)=u_{0}\left(x\right)\in\mathcal{D}(\mathbb{Z}_{p}).\end{array}\right$.
By passing to the Fourier transform in (4.1),
$\left\{\begin{array}[c]{l}\frac{\partial}{\partial t}\widehat{u}\left(\xi,t\right)=\left(\widehat{J}(\left|\xi\right|_{p})-1\right)\widehat{u}\left(\xi,t\right)\text{, }\xi\in\mathbb{Q}_{p},t\in\left[0,T\right]\\ \\ \widehat{u}\left(\xi,0\right)=\widehat{u}_{0}\left(\xi\right)\in\mathcal{D}(\mathbb{Q}_{p}).\end{array}\right$.
Then, $\widehat{\text{ }u}\left(\xi,t\right)=\mathrm{e}^{t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\widehat{u}_{0}\left(\xi\right)$, and setting $Z_{0}(x,t)=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\mathrm{e}^{t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\right)$, and taking formally the inverse Fourier transform,
$u\left(x,t\right)$ $=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\mathrm{e}^{t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\widehat{u}_{0}\left(\xi\right)\right)=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\mathrm{e}^{t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\right)\ast u_{0}\left(x\right)$
$=Z_{0}(x,t)\ast u_{0}\left(x\right).$
In Appendix B, using [40, Theorem 3.1], we show that $u\left(x,t\right)=Z_{0}\left(x,t\right)\ast u_{0}\left(x\right)$ is the unique solution of (4.1), in a suitable function space, and that
$p(t,x,B)=\begin{cases}\int_{B}Z_{0}(x-y,t)dy=1_{B}\left(x\right)\ast Z_{0}(x,t)&\text{for }t>0,\quad x\in\mathbb{Z}_{p},\\ 1_{B}(x)&\text{for }t=0,\end{cases}$
is the transition function of a Markov process. Furthermore,
${\displaystyle\int\limits_{\mathbb{Z}_{p}}}Z_{0}(x,t)dx=1\text{, for }t>0,$
see Theorem 9.1, in Appendix B.
4.2. CTMCs
In this section, we discuss the results given in Appendix C. We take $\mathcal{K}_{r}$, $r\in\mathbb{J}$ as in (3.6), where $\mathbb{J}$ is a countable set, possibly infinite. The value
$p_{r,v}(t):={\displaystyle\int\limits_{\mathcal{K}_{v}}}p(t,x,\mathcal{K}_{r})\frac{dx}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)},$
with $\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)=\int_{\mathcal{K}_{v}}dx$, gives the transition probability, per unit of time, that a particle/walker starting at some position in $\mathcal{K}_{v}$ will be found in the set $\mathcal{K}_{r}$ at the time $t$. Furthermore, $\sum_{r\in\mathbb{J}}p_{r,v}(t)=1$, for $t>0$, cf. Lemma 10.1 in Appendix C.
Now, $\left[p_{r,v}(t)\right]_{r,v\in\mathbb{J}}$ is the transition matrix of a CTMC. Furthermore, under a suitable hypothesis on the Fourier transform of kernel $J(\left|x\right|_{p})$, the CTMC admits a stationary distribution $\left[p_{r}^{\text{sta}}\right]_{r\in\mathbb{J}}=\left[\int_{\mathcal{K}_{r}}dx\right]_{r\in\mathbb{J}}$; see Appendix C, Theorem 10.1. In Lemma 12.2, in Appendix E, we compute the rate matrix for Markov chains with finite number of states. The existence of a stationary distribution for this type of CTMC follows from well-known results. Here, we work with CTMCs with countable states (finite or infinite) which are obtained as discretizations of $p$-adic heat equations; see [40], and the references therein.
5. $p$-adic Schrödinger equations and CTQMCs
5.1. A class of $p$-adic Schrödinger equations
By performing a Wick rotation $t\rightarrow\mathrm{i}t$ in (4.1), and taking $\Psi\left(x,t\right)=u\left(x,it\right)$, we obtain the following Schrödinger equations in the unit ball:
(5.1) $\left\{\begin{array}[c]{l}\mathrm{i}\frac{\partial}{\partial t}\Psi\left(x,t\right)=-J\left(\left|x\right|_{p}\right)\ast\Psi\left(x,t\right)+\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\geq 0\\ \\ \Psi\left(x,0\right)=\psi_{0}\left(x\right)\in\mathcal{D}(\mathbb{Z}_{p}).\end{array}\right$.
By using the Fourier transform, we obtain that the solution of (5.1) is given by
(5.2) $\Psi\left(x,t\right)=Z_{0}(x,\mathrm{i}t)\ast\psi_{0}\left(x\right)\text{ in }\mathcal{D}^{\prime}(\mathbb{Q}_{p})\text{,}$
cf. Appendix D. This result is useful for calculations; however, it is necessary to show that if $\psi_{0}(x)\in L^{2}\left(\mathbb{Z}_{p}\right)$, with $\left\|\psi_{0}\right\|_{2}=1$, then $\left\|\Psi(\cdot,t)\right\|_{2}=1$ for any $t>0$. This fact is established in Theorem 11.1, in Appendix D.
5.2. CTQMCs
We use Theorem 3.1 to construct a family of CTQMCs: for $l\geq 1$, set $\mathbb{J}=G_{l}$ and
(5.3) $\pi_{r,v}\left(t\right)=\left|\left\langle p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right),Z_{0}(x,it)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\right\rangle\right|^{2},$
for $r,v\in G_{l}$. Then, $\left[\pi_{r,v}\left(t\right)\right]_{r,v\in G_{l}}$ is the transition matrix of a CTQMC. Furthermore, $\pi_{r,v}\left(t\right)$ is an oscillatory integral:
$\pi_{r,v}\left(t\right)=p^{-2l}\left|\text{ }{\displaystyle\int\limits_{\left|z\right|_{p}\leq p^{l}}}\chi_{p}\left(z\left(v-r\right)\right)e^{-it\widehat{J}\left(\left|z\right|_{p}\right)}dz\right|^{2};$
see Proposition 11.1, Theorem 11.2, in Appendix D.
6. Discretizations of $p$-adic evolutionary differential equations
The numerical simulations required here use spatial discretizations of the $p$-adic heat and Schrödinger equations. An in-depth discussion is given in Appendix D.
For $l\geq 1$, the $p$-adic integers of the form
(6.1) $I=I_{0}+I_{1}p+\ldots+I_{l-1}p^{l-1},$
with $I_{j}\in\left\{0,\ldots,p-1\right\}$, form a set of representatives of the elements of the quotient group $G_{l}=\mathbb{Z}_{p}/p^{l}\mathbb{Z}_{p}$. From a geometric perspective, $G_{l}$ corresponds to a finite rooted tree, with $l+1$ levels; the vertices at the top level correspond with the $p$-adic integers of the form (6.1). We denote by $\Omega\left(p^{l}\left|x-I\right|_{p}\right)$ the characteristic function of the ball $I+p^{l}\mathbb{Z}_{p}$. We denote by $\mathcal{D}_{l}(\mathbb{Z}_{p})$, the $\mathbb{C}$-vector space spanned by the linearly independent set $\left\{\Omega\left(p^{l}\left|x-I\right|_{p}\right)\right\}_{I\in G_{l}}$. The space of test functions $\mathcal{D}(\mathbb{Z}_{p})$ satisfies
$\mathcal{D}(\mathbb{Z}_{p})={\displaystyle\bigcup\limits_{l=0}^{\infty}}\mathcal{D}_{l}(\mathbb{Z}_{p})\text{, with }\mathcal{D}_{0}(\mathbb{Z}_{p})=\mathbb{C}\text{, and }\mathcal{D}_{l}(\mathbb{Z}_{p})\subset\mathcal{D}_{l+1}(\mathbb{Z}_{p})\text{.}$
Furthermore, since the Haar measure of $\mathbb{Z}_{p}$ is one, we have
$L^{1}(\mathbb{Z}_{p})\supset L^{\rho}(\mathbb{Z}_{p})\supset\mathcal{C}(\mathbb{Z}_{p})\supset\mathcal{D}(\mathbb{Z}_{p}),\text{ for }\rho\in\left(1,\infty\right]\text{,}$
and $\mathcal{D}(\mathbb{Z}_{p})$ is dense in $L^{1}(\mathbb{Z}_{p})$, which implies that a function $f$ $\in$ $L^{1}(\mathbb{Z}_{p})$ can be approximated by a function $\varphi\in$ $\mathcal{D}_{l}(\mathbb{Z}_{p})$, with $l=l(f)$. This fact implies the existence of a linear bounded operator
$\boldsymbol{P}_{l}:L^{1}(\mathbb{Z}_{p})\rightarrow\mathcal{D}_{l}(\mathbb{Z}_{p})$
such that $\boldsymbol{P}_{l}f\in\mathcal{D}_{l}(\mathbb{Z}_{p})$ is an approximation of $f$ $\in$ $L^{1}(\mathbb{Z}_{p})$, see Appendix E for details.
If $\boldsymbol{G}:L^{1}(\mathbb{Z}_{p})\rightarrow L^{1}(\mathbb{Z}_{p})$ is a linear operator, a discretization $\boldsymbol{G}^{\left(l\right)}:\mathcal{D}_{l}(\mathbb{Z}_{p})\rightarrow\mathcal{D}_{l}(\mathbb{Z}_{p})$ is constructed as $\boldsymbol{G}^{\left(l\right)}=\boldsymbol{P}_{l}\boldsymbol{G}$. Since $\mathcal{D}_{l}(\mathbb{Z}_{p})$ is a finite-dimensional vector space, $\boldsymbol{G}^{\left(l\right)}$ is just a matrix.
On the other hand, the space $\mathcal{D}_{l}(\mathbb{Z}_{p})$ is invariant under the operator $\boldsymbol{J}f=J\ast f-f$, cf. Lemma 12.2 in Appendix E; in this way, we construct spatial approximations of the $p$-adic heat and Schrödinger equations (4.1)-(5.1); see equations (12.5)-(12.6) in Appendix E. We compute explicitly the matrix $\boldsymbol{J}^{\left(l\right)}$, see Lemma 12.2 in Appendix E.
We identify $\mathcal{D}_{l}(\mathbb{Z}_{p})$ with the Hilbert space $\mathbb{C}^{p^{l}}$. This identification maps
$\left\{p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)\right\}_{r\in G_{l}}$
into the standard orthonormal basis $\left\{e_{r}\right\}$ $_{r\in G_{l}}$ of $\mathbb{C}^{p^{l}}$. We show that the transition probabilities of the CTQMC, $\left[\pi_{r,v}\left(t\right)\right]_{r,v\in G_{l}}$, given in (5.3) satisfy $\pi_{I,J}\left(t\right)=\left|\left\langle e_{I}\right|\mathrm{e}^{it\boldsymbol{J}^{\left(l\right)}}\left|e_{J}\right\rangle_{l}\right|^{2}$, for $J,I\in G_{l}$. We use this formula in the numerical simulations given in the next section.
7. Numerical simulations
We want to compare (via numerical simulation) the CTQMCs against a classical CTMCs. We take $r\in\mathbb{J}=G_{l}$, and $\mathcal{K}_{r}=B_{-l}(r)=r+p^{l}\mathbb{Z}_{p}$, and identify $\left\{p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)\right\}_{r\in G_{l}}$ with the standard orthonormal basis $\left\{e_{r}\right\}_{r\in G_{l}}$ of $\mathbb{R}^{p^{l}}$. Then
$p_{r,v}(t)$ $=$ $p^{l}\int\limits_{B_{-l}(v)}Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)dx$
$=$ $\left\langle\Omega\left(p^{l}\left|x-v\right|_{p}\right),Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)\right\rangle,$
for $r,v\in G_{l}$. Now, $Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)$ is the solution of the Cauchy problem for the heat equation with initial datum $\Omega\left(p^{l}\left|x-r\right|_{p}\right)$, then by Remark 12.4,
$Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)=\mathrm{e}^{-t\boldsymbol{J}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)=\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}e_{r},$
and
(7.1) $p_{r,v}(t)=\left\langle e_{v},\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}e_{r}\right\rangle_{l}.$
7.1. Heat and Schrödinger equation attached to Bessel potentials
Set $\Gamma\left(\alpha\right)=\frac{1-p^{\alpha-1}}{1-p^{-\alpha}}$, $\alpha\in\mathbb{R}\smallsetminus\left\{0\right\}$. We set for $\alpha\neq 1$,
$J_{\alpha}\left(x\right)=J_{\alpha}\left(\left|x\right|_{p}\right)=\frac{1}{\Gamma\left(\alpha\right)}\left\{\left|x\right|_{p}^{\alpha-1}-p^{\alpha-1}\right\}\Omega\left(\left|x\right|_{p}\right),$
and
$J_{1}\left(x\right)=J_{1}\left(\left|x\right|_{p}\right)=\left(1-p^{-1}\right)\log_{p}\left(\frac{p}{\left|x\right|_{p}}\right)\Omega\left(\left|x\right|_{p}\right).$
Notice that for $\alpha\in\left(0,1\right]$, $J_{\alpha}\left(\left|x\right|_{p}\right)$ is not defined at the origin.
Then $\left\|J_{\alpha}\right\|_{1}=1$, $J_{\alpha}\left(|\xi|_{p}\right)\geq 0$, for $x\neq 0$, and $\widehat{J}_{\alpha}\left(|\xi|_{p}\right)=\left(\max\left\{1,\left|\xi\right|_{p}\right\}\right)^{-\alpha}$, see [41, Lemma 5.2].
For $\alpha\in\left(0,\infty\right)$, we define the operator
$\boldsymbol{J}_{\alpha}\varphi\left(x\right)=J_{\alpha}\left(\left|x\right|_{p}\right)\ast\varphi\left(x\right)-\varphi\left(x\right)\text{, for }\varphi\in\mathcal{C}(\mathbb{Z}_{p}),$
the heat equation
$\frac{\partial}{\partial t}u\left(x,t\right)=J_{\alpha}\left(\left|x\right|_{p}\right)\ast u\left(x,t\right)-u\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\geq 0,$
and the Schrödinger equation
$\mathrm{i}\frac{\partial}{\partial t}\Psi\left(x,t\right)=-\left\{J_{\alpha}\left(\left|x\right|_{p}\right)\ast\Psi\left(x,t\right)-\Psi\left(x,t\right)\right\}\text{, }x\in\mathbb{Z}_{p},t\geq 0.$
7.2. Numerical simulation 1
In this section, for $\alpha\in\left(0,\infty\right)\smallsetminus\left\{1\right\}$, we provide numerical simulations for the CTMCs and CTQMCs attached to the heat and Schrödinger equations attached to Bessel potentials.
Matrix of operator $\boldsymbol{J}_{\alpha}$
(7.2) $\boldsymbol{J}^{\left(l\right)}\left(\alpha\right)=\left[J_{I,K}^{\left(l\right)}\left(\alpha\right)\right]_{I,K\in G_{l}},$
(7.3) $J_{I,K}^{\left(l\right)}\left(\alpha\right)=\left\{\begin{array}[c]{lll}p^{-l}\frac{1-p^{-\alpha}}{1-p^{\alpha-1}}\left\{\left|I-K\right|_{p}^{\alpha-1}-p^{\alpha-1}\right\}&\text{if}&I\neq K\\ &&\\ -{\displaystyle\sum\limits_{\begin{subarray}{c}I\in G_{l}\\ I\neq 0\end{subarray}}}p^{-l}\frac{1-p^{-\alpha}}{1-p^{\alpha-1}}\left\{\left|I\right|_{p}^{\alpha-1}-p^{\alpha-1}\right\}&\text{if}&I=K\end{array}\right$.
Figure 1. Numerical Simulation 1. Matrix $\boldsymbol{J}^{\left(l\right)}\left(\alpha\right)$. The parameters are $p=2$, $l=5$, $\alpha=1.2$. The CTMC has $32$ states organized in a finite tree $G_{5}$. The figure illustrates the ultrametric nature of matrix $\boldsymbol{J}^{\left(l\right)}\left(\alpha\right)$.
CTMC
$p_{I,J}(t)=\left\langle e_{J},\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}\left(\alpha\right)}e_{I}\right\rangle_{l}$
Figure 2. Numerical Simulation 1. $p_{I,12}(t)$ versus $\pi_{I,12}\left(t\right)$. The state II runs through 55 states. The parameters are $p=2$, $l=5$, $\alpha=1.2$. The figure (a) shows that $\lim_{t\rightarrow\infty}p_{I,12}(t)\approx 2^{-5}$, while figure (b) illustrates that $\lim_{t\rightarrow\infty}\pi_{I,12}\left(t\right)$ does not exist.
CTQMC
$\pi_{I,J}\left(t\right)=\left|\left\langle e_{I}\right|\mathrm{e}^{\mathrm{i}t\boldsymbol{J}^{\left(l\right)}\left(\alpha\right)}\left|e_{J}\right\rangle_{l}\right|^{2}\text{, for }J,I\in G_{l},$
for $\alpha\in\left(0,\infty\right)\smallsetminus\left\{1\right\}$. Here, it is relevant to mention, that $\boldsymbol{J}^{\left(l\right)}\left(\alpha\right)$ is a symmetric matrix, and thus $\pi_{I,J}\left(t\right)$ is well-defined for $\alpha\in\left(0,\infty\right)\smallsetminus\left\{1\right\}$.
Figure 3. Numerical Simulation 1. $\pi_{I,J}\left(t\right)$ for six different times: t=0t=0,11, $200$, $500$, $1000$, $4000$, $10000$. The states $I$, $J$ run through $32$ states. The parameters are $p=2$, $l=5$, $\alpha=1.2$.
Figure 4. Numerical Simulation 1. $\pi_{I,J}\left(200\right)$ for six different values of the parameter $\alpha$. The other parameters are $p=2$, $l=5$. The states $I$, $J$ run through $32$ states. The values $\alpha=0$, 11 correspond to a pole, respectively a zero, of $\Gamma\left(\alpha\right)$. When the value of $\alpha$ is near to 0, 11, or $\alpha\geq 5$, the transitions of the CTQMC occur around the diagonal; the walker only performs very short walks around each state. While for other values of $\alpha$, the walker performs long walks around any state.
Figure 5. Numerical Simulation 1. The limiting distribution (or longtime average) is defined as
$\chi_{I,J}=\lim_{T\rightarrow\infty}\frac{1}{T}{\displaystyle\int\limits_{0}^{T}}\pi_{I,J}\left(t\right)dt.$
A numerical approximation of this limiting distribution, $p=2$, $l=5$, $\alpha=1.2$ is shown in the figure. The states $I$, $J$ run through $32$ states. In the literature, this limiting distribution is compared with
$p_{I}^{\text{sta}}=\lim_{t\rightarrow\infty}p_{I,J}\left(t\right)=p^{-l},$
cf. Theorem 10.1, in Appendix C. The simulation shows that $\chi_{I,J}>p_{I}^{\text{sta}}=2^{-5}$. This fact is interpreted as the computational power of the CTQMC is greater than that of the corresponding CTMC. The computation of the average used the time interval $\left[0,10000\right]$.
7.3. Numerical simulation 2
In this section, for $\alpha=1$, we provide numerical simulations for the CTMCs and CTQMCs attached to the heat and Schrödinger equations attached to Bessel potentials.
Matrix of operator ${\boldsymbol{J}}_{1}$
${\boldsymbol{J}}^{\left(l\right)}\left(1\right)=\left[J_{I,K}^{\left(l\right)}\left(1\right)\right]_{I,K\in G_{l}},\text{ }$
$J_{I,K}^{\left(l\right)}\left(1\right)=\left\{\begin{array}[c]{lll}p^{-l}\left(1-p^{-1}\right)\log_{p}\left(\frac{p}{\left|I-K\right|_{p}}\right)&\text{if}&I\neq K\\ &&\\ -{\displaystyle\sum\limits_{\begin{subarray}{c}I\in G_{l}\\ I\neq 0\end{subarray}}}p^{-l}\left(1-p^{-1}\right)\log_{p}\left(\frac{p}{\left|I\right|_{p}}\right)&\text{if}&I=K.\end{array}\right$.
CTMC
$p_{r,v}(t)=\left\langle e_{v},\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}\left(1\right)}e_{r}\right\rangle_{l}.$
CTQMC
$\pi_{I,J}\left(t\right)=\left|\left\langle e_{I}\right|\mathrm{e}^{\mathrm{i}t\boldsymbol{J}^{\left(l\right)}\left(1\right)}\left|e_{J}\right\rangle_{l}\right|^{2}\text{, for }J,I\in G_{l}.$
Figure 6. Numerical Simulation 2. $p_{I,26}(t)$ versus $\pi_{I,26}(t)$ for 88 different states. The parameters are $p=2$, $l=5$, $\alpha=1$. The figure (a) clearly illustrates that $\lim_{t\rightarrow\infty}p_{I,26}(t)\approx 2^{-5}$.
Figure 7. Numerical Simulation 2. $\pi_{I,J}\left(t\right)$ for six different times: t=0t=0,11, $200$, $500$, $1000$, $4000$, $10000$. The states $I$, $J$ run through $32$ states. The parameters are $p=2$, $l=5$, $\alpha=1$. Notice that the figures 12,15, 21 in [3] also show an ultrametric nature.
Figure 8. Numerical Simulation 2. A numerical approximation of the limiting distribution $\chi_{I,J}$. The computation of the average used the time interval $\left[0,10000\right]$. The parameters are $p=2$, $l=5$, $\alpha=1$.
Figure 9. Limiting distribution $\chi=$ $\left[\chi_{I,J}\right]$ for several values of $\alpha$. The parameters are $p=2$, $l=5$. The computation of the average used the time interval $\left[0,10000\right]$. This figure illustrates the strong influence of $\alpha$ on the limiting distribution. For $\alpha\geq 4.5$, the distribution $\chi$ is concentrated around the diagonal, which means that only very short walks around each state are allowed.
Figure 10. For $\chi=$ $\left[\chi_{I,J}\right]$, we set $\left\|\chi\right\|_{\text{max}}=\max_{I,J}\left\{\chi_{I,J}\right\}$. The figure shows $\left\|\chi\right\|_{\text{max}}$ for several values of $\alpha$. The other parameters are $p=2$, $l=5$. The computation of the average used the time interval $\left[0,10000\right]$. This figure shows an alternative way to visualize the information displayed in Figure 9.
8. Appendix A: Basic facts on $p$-adic analysis
In this section, we fix the notation and collect some basic results on $p$-adic analysis that we will use throughout the article. For a detailed exposition on $p$-adic analysis, the reader may consult [19], [32], [35], [41].
8.1. The field of $p$-adic numbers
The field of $p$-adic numbers $\mathbb{Q}_{p}$ is defined as the completion of the field of rational numbers $\mathbb{Q}$ with respect to the $p$-adic norm $|\cdot|_{p}$, which is defined as
$|x|_{p}=\begin{cases}0&\text{if }x=0\\ p^{-\gamma}&\text{if }x=p^{\gamma}\dfrac{a}{b},\end{cases}$
where $a$ and $b$ are integers coprime with $p$. The integer $\gamma=ord_{p}(x):=ord(x)$, with $ord(0):=+\infty$, is called the $p$-adic order of $x$.
The metric space $\left(\mathbb{Q}_{p},|\cdot|_{p}\right)$ is a complete ultrametric space. As a topological space $\mathbb{Q}_{p}$ is homeomorphic to a Cantor-like subset of the real line, see, e.g., [19], [35].
Any $p$-adic number $x\neq 0$ has a unique expansion of the form
$x=p^{ord(x)}\sum_{j=0}^{\infty}x_{j}p^{j},$
where $x_{j}\in\{0,1,2,\dots,p-1\}$ and $x_{0}\neq 0$. By using this expansion, we define the fractional part $\{x\}_{p}$ of $x\in\mathbb{Q}_{p}$ as the rational number
$\{x\}_{p}=\begin{cases}0&\text{if }x=0\text{ or }ord(x)\geq 0\\ p^{ord(x)}\sum_{j=0}^{-ord(x)-1}x_{j}p^{j}&\text{if }ord(x)<0.\end{cases}$
In addition, any $x\in\mathbb{Q}_{p}\smallsetminus\left\{0\right\}$ can be represented uniquely as $x=p^{ord(x)}v$, where $\left|v\right|_{p}=1$.
8.2. Topology of $\mathbb{Q}_{p}$
For $r\in\mathbb{Z}$, denote by $B_{r}(a)=\{x\in\mathbb{Q}_{p};\left|x-a\right|_{p}\leq p^{r}\}$ the ball of radius $p^{r}$ with center at $a\in\mathbb{Q}_{p}$, and take $B_{r}(0):=B_{r}$. Notice that $B_{0}=\mathbb{Z}_{p}$ is the ring of $p$-adic integers. We also denote by $S_{r}(a)=\{x\in\mathbb{Q}_{p};\left|x-a\right|_{p}=p^{r}\}$ the sphere of radius $p^{r}$ with center at $a\in\mathbb{Q}_{p}$, and take $S_{r}(0):=S_{r}$. We notice that $S_{0}=\mathbb{Z}_{p}^{\times}$ (the group of units of $\mathbb{Z}_{p}$). The balls and spheres are both open and closed subsets in $\mathbb{Q}_{p}$. In addition, two balls in $\mathbb{Q}_{p}$ are either disjoint or one is contained in the other.
As a topological space $\left(\mathbb{Q}_{p},\left|\cdot\right|_{p}\right)$ is totally disconnected, i.e., the only connected subsets of $\mathbb{Q}_{p}$ are the empty set and the points. A subset of $\mathbb{Q}_{p}$ is compact if and only if it is closed and bounded in $\mathbb{Q}_{p}$, see, e.g., [19, Section 1.3], or [35, Section 1.8]. The balls and spheres are compact subsets. Thus $\left(\mathbb{Q}_{p},\left|\cdot\right|_{p}\right)$ is a locally compact topological space.
8.3. The Haar measure
Since $(\mathbb{Q}_{p},+)$ is a locally compact topological group, there exists a Haar measure $dx$, which is invariant under translations, i.e., $d(x+a)=dx$, [42]. If we normalize this measure by the condition $\int_{\mathbb{Z}_{p}}dx=1$, then $dx$ is unique.
Notation 1.
We will use $\Omega\left(p^{-r}\left|x-a\right|_{p}\right)$ to denote the characteristic function of the ball $B_{r}(a)=a+p^{-r}\mathbb{Z}_{p}$, where
$\mathbb{Z}_{p}=\left\{x\in\mathbb{Q}_{p};\left|x\right|_{p}\leq 1\right\}$
is the $1$-dimensional unit ball. For more general sets, we will use the notation $1_{A}$ for the characteristic function of set $A$.
8.4. The Bruhat-Schwartz space
A complex-valued function $\varphi$ defined on $\mathbb{Q}_{p}$ is called locally constant if for any $x\in\mathbb{Q}_{p}$ there exist an integer $l(x)\in\mathbb{Z}$ such that
(8.1) $\varphi(x+x^{\prime})=\varphi(x)\text{ for any }x^{\prime}\in B_{l(x)}.$
A function $\varphi:\mathbb{Q}_{p}\rightarrow\mathbb{C}$ is called a Bruhat-Schwartz function (or a test function) if it is locally constant with compact support. Any test function can be represented as a linear combination, with complex coefficients, of characteristic functions of balls. The $\mathbb{C}$-vector space of Bruhat-Schwartz functions is denoted by $\mathcal{D}(\mathbb{Q}_{p})$. For $\varphi\in\mathcal{D}(\mathbb{Q}_{p})$, the largest number $l=l(\varphi)$ satisfying (8.1) is called the exponent of local constancy (or the parameter of constancy) of $\varphi$.
8.5. $L^{\rho}$ spaces
Given $\rho\in[1,\infty)$, we denote by $L^{\rho}\left(\mathbb{Q}_{p}\right)$ the $\mathbb{C}$-vector space of all the complex valued functions $g$ satisfying
$\left\|g\right\|_{\rho}=\left(\text{ }{\displaystyle\int\limits_{\mathbb{Q}_{p}}}\left|g\left(x\right)\right|^{\rho}dx\right)^{\frac{1}{\rho}}<\infty,$
where $dx$ is the normalized Haar measure on $\left(\mathbb{Q}_{p},+\right)$.
If $U$ is an open subset of $\mathbb{Q}_{p}$, $\mathcal{D}(U)$ denotes the $\mathbb{C}$-vector space of test functions with supports contained in $U$, then $\mathcal{D}(U)$ is dense in
$L^{\rho}\left(U\right)=\left\{\varphi:U\rightarrow\mathbb{C};\left\|\varphi\right\|_{\rho}=\left\{{\displaystyle\int\limits_{U}}\left|\varphi\left(x\right)\right|^{\rho}dx\right\}^{\frac{1}{\rho}}<\infty\right\},$
for $1\leq\rho<\infty$, see, e.g., [35, Section 4.3].
8.6. The Fourier transform
Set $\chi_{p}(y)=\exp(2\pi i\{y\}_{p})$ for $y\in\mathbb{Q}_{p}$. The map $\chi_{p}(\cdot)$ is an additive character on $\mathbb{Q}_{p}$, i.e., a continuous map from $\left(\mathbb{Q}_{p},+\right)$ into $S$ (the unit circle considered as a multiplicative group) satisfying $\chi_{p}(x_{0}+x_{1})=\chi_{p}(x_{0})\chi_{p}(x_{1})$, $x_{0},x_{1}\in\mathbb{Q}_{p}$. The additive characters of $\mathbb{Q}_{p}$ form an Abelian group which is isomorphic to $\left(\mathbb{Q}_{p},+\right)$. The isomorphism is given by $\kappa\rightarrow\chi_{p}(\kappa x)$, see, e.g., [35, Section 2.3].
The Fourier transform of $\varphi\in\mathcal{D}(\mathbb{Q}_{p})$ is defined as
$\mathcal{F}\varphi(\xi)={\displaystyle\int\limits_{\mathbb{Q}_{p}}}\chi_{p}(\xi x)\varphi(x)dx\quad\text{for }\xi\in\mathbb{Q}_{p}.$
The Fourier transform is a linear isomorphism from $\mathcal{D}(\mathbb{Q}_{p})$ onto itself satisfying
(8.2) $(\mathcal{F}(\mathcal{F}\varphi))(\xi)=\varphi(-\xi),$
see, e.g., [35, Section 4.8]. We will also use the notation $\mathcal{F}_{x\rightarrow\kappa}\varphi$ and $\widehat{\varphi}$ for the Fourier transform of $\varphi$.
The Fourier transform extends to $L^{2}$. If $f\in L^{2}\left(\mathbb{Q}_{p}\right)$, its Fourier transform is defined as
[…]
8.7. Distributions
The $\mathbb{C}$-vector space $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$ of all continuous linear functionals on $\mathcal{D}(\mathbb{Q}_{p})$ is called the Bruhat-Schwartz space of distributions. Every linear functional on $\mathcal{D}(\mathbb{Q}_{p})$ is continuous, i.e., $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$ agrees with the algebraic dual of $\mathcal{D}(\mathbb{Q}_{p})$, see, e.g., [19, Chapter 1, VI.3, Lemma].
We endow $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$ with the weak topology, i.e. a sequence $\left\{T_{j}\right\}_{j\in\mathbb{N}}$ in $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$ converges to $T$ if $\lim_{j\rightarrow\infty}T_{j}\left(\varphi\right)=T\left(\varphi\right)$ for any $\varphi\in\mathcal{D}(\mathbb{Q}_{p})$. The map
$\begin{array}[c]{lll}\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)\times\mathcal{D}(\mathbb{Q}_{p})&\rightarrow&\mathbb{C}\\ \left(T,\varphi\right)&\rightarrow&T\left(\varphi\right)\end{array}$
is a bilinear form which is continuous in $T$ and $\varphi$ separately. We call this map the pairing between $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$ and $\mathcal{D}(\mathbb{Q}_{p})$. From now on we will use $\left(T,\varphi\right)$ instead of $T\left(\varphi\right)$.
8.8. The Fourier transform of a distribution
The Fourier transform $\mathcal{F}\left[T\right]$ of a distribution $T\in\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$ is defined by
$\left(\mathcal{F}\left[T\right],\varphi\right)=\left(T,\mathcal{F}\left[\varphi\right]\right)\text{ for all }\varphi\in\mathcal{D}\left(\mathbb{Q}_{p}\right)\text{.}$
The Fourier transform $T\rightarrow\mathcal{F}\left[T\right]$ is a linear and continuous isomorphism from $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$ onto $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$. Furthermore, $T=\mathcal{F}\left[\mathcal{F}\left[T\right]\left(-\xi\right)\right]$.
Let $T\in\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$ be a distribution. Then supp $T\subset B_{L}$ if and only if $\mathcal{F}\left[T\right]$ is a locally constant function, and the exponent of local constancy of $\mathcal{F}\left[T\right]$ is $\geq-L$. In addition
$\mathcal{F}\left[T\right]\left(\xi\right)=\left(T\left(y\right),\Omega\left(p^{-L}\left|y\right|_{p}\right)\chi_{p}\left(\xi y\right)\right),$
see, e.g., [35, Section 4.9].
8.9. The direct product of distributions
Given $F,G\in\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$, their direct product $F\times G$ is defined by the formula
$\left(F\left(x\right)\times G\left(y\right),\varphi\left(x,y\right)\right)=\left(F\left(x\right),\left(G\left(y\right),\varphi\left(x,y\right)\right)\right)\text{ for }\varphi\left(x,y\right)\in\mathcal{D}\left(\mathbb{Q}_{p}^{2}\right).$
The direct product is commutative: $F\times G=G\times F$. In addition the direct product is continuous with respect to the joint factors.
8.10. The convolution of distributions
Given $F,G\in\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$, their convolution $F\ast G$ is defined by
$\left(F\ast G,\varphi\right)=\lim_{k\rightarrow\infty}\left(F\left(y\right)\times G\left(x\right),\Omega\left(p^{-k}\left|y\right|_{p}\right)\varphi\left(x+y\right)\right)$
[…]
$\left(F\ast G,\varphi\right)=\left(F\left(y\right)\times G\left(x\right),\Omega\left(p^{-L}\left|y\right|_{p}\right)\varphi\left(x+y\right)\right)\text{ for }\varphi\in\mathcal{D}\left(\mathbb{Q}_{p}\right).$
In the case in which $G=\psi\in\mathcal{D}\left(\mathbb{Q}_{p}\right)$, $F\ast\psi$ is a locally constant function given by
$\left(F\ast\psi\right)\left(y\right)=\left(F\left(x\right),\psi\left(y-x\right)\right),$
see, e.g., [19, Section 7.1].
8.11. The multiplication of distributions
Set $\delta_{k}\left(x\right):=p^{k}\Omega\left(p^{k}\left|x\right|_{p}\right)$ for $k\in\mathbb{N}$. Given $F,G\in\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$, their product $F\cdot G$ is defined by
$\left(F\cdot G,\varphi\right)=\lim_{k\rightarrow\infty}\left(G,\left(F\ast\delta_{k}\right)\varphi\right)$
if the limit exists for all $\varphi\in\mathcal{D}\left(\mathbb{Q}_{p}\right)$. If the product $F\cdot G$ exists then the product $G\cdot F$ exists and they are equal.
We recall that the existence of the product $F\cdot G$ is equivalent to the existence of $\mathcal{F}\left[F\right]\ast\mathcal{F}\left[G\right]$. In addition, $\mathcal{F}\left[F\cdot G\right]=\mathcal{F}\left[F\right]\ast\mathcal{F}\left[G\right]$ and $\mathcal{F}\left[F\ast G\right]=\mathcal{F}\left[F\right]\cdot\mathcal{F}\left[G\right]$, see, e.g., [19, Section 7.5].
9. Appendix B: $p$-adic heat equations on the unit ball and Markov processes
9.1. Preliminary results
For fix a real-valued function $J:\left[0,1\right]\rightarrow\left[0,\infty\right)$, we define
$\begin{array}[c]{cccc}J:&\mathbb{Z}_{p}&\rightarrow&\left[0,\infty\right)\\ &x&\rightarrow&J(\left|x\right|_{p}).\end{array}$
We assume that
$\left\|J\right\|_{L^{1}\left(\mathbb{Z}_{p}\right)}:=\left\|J(\left|x\right|_{p})\right\|_{1}={\displaystyle\int\limits_{\mathbb{Z}_{p}}}J(\left|x\right|_{p})dx=1.$
Remark 9.1.
(i) By extending $J(\left|x\right|_{p})$ as zero outside of the unit ball, we have $J(\left|x\right|_{p})\in L^{1}\left(\mathbb{Q}_{p}\right)$, and then its Fourier transform $\widehat{J}(\xi):\mathbb{Q}_{p}\rightarrow\mathbb{C}$ is well-defined. The Fourier transform does not preserve the support of $J(\left|x\right|_{p})$, i.e. $\widehat{J}(\xi)\neq 0$ for points $\xi$ outside of the unit ball.
(ii) Notice that $\widehat{J}(0)=\left\|J(\left|x\right|_{p})\right\|_{1}=1$.
Lemma 9.1.
With the above notation, the following assertions hold:
(i) the Fourier transform $\widehat{J}(\xi)$ of $J(\left|x\right|_{p})$ is a real-valued function given by
$\widehat{J}(\xi)={\displaystyle\sum\limits_{j=0}^{\infty}}J(p^{-j})\left\{p^{-j}\Omega\left(p^{-j}\left|\xi\right|_{p}\right)-p^{-j-1}\Omega\left(p^{-j-1}\left|\xi\right|_{p}\right)\right\}\text{, }\xi\in\mathbb{Q}_{p};$
(ii) for $\xi\in\mathbb{Q}_{p}\smallsetminus\left\{0\right\}$, $\widehat{J}(\xi)=\widehat{J}(u\xi)$, for any $u\in\mathbb{Z}_{p}^{\times}$, this means that $\widehat{J}(\xi)$ is a radial function $\widehat{J}(\xi)=\widehat{J}(\left|\xi\right|_{p})$ for $\xi\in\mathbb{Q}_{p}\smallsetminus\left\{0\right\}$;
(iii) for $\xi\in\mathbb{Q}_{p}$, $\widehat{J}(\xi)=\widehat{J}(\xi+\xi_{0})$ for any $\xi\in\mathbb{Z}_{p}$. In particular $\widehat{J}(0)=\widehat{J}(\xi_{0})=\left\|J(\left|x\right|_{p})\right\|_{1}$, for $\xi_{0}\in\mathbb{Z}_{p}$. Then $\widehat{J}(\xi)=\widehat{J}(\left|\xi\right|_{p})$ for $\xi\in\mathbb{Q}_{p}$.
Proof.
(i) By using the partition,
$\mathbb{Z}_{p}\smallsetminus\left\{0\right\}={\displaystyle\bigsqcup\limits_{j=0}^{\infty}}p^{j}\mathbb{Z}_{p}^{\times},$
and the fact that $\mu_{\text{Haar}}\left(\left\{0\right\}\right)=0$,
$\widehat{J}(\xi)$ $={\displaystyle\int\limits_{\mathbb{Z}_{p}}}J(\left|x\right|_{p})\chi_{p}\left(\xi x\right)dx={\displaystyle\int\limits_{\mathbb{Z}_{p}\smallsetminus\left\{0\right\}}}J(\left|x\right|_{p})\chi_{p}\left(\xi x\right)dx={\displaystyle\sum\limits_{j=0}^{\infty}}\text{ }{\displaystyle\int\limits_{p^{j}\mathbb{Z}_{p}^{\times}}}J(\left|x\right|_{p})\chi_{p}\left(\xi x\right)dx$
(9.1) $={\displaystyle\sum\limits_{j=0}^{\infty}}J(p^{-j}){\displaystyle\int\limits_{p^{j}\mathbb{Z}_{p}^{\times}}}\chi_{p}\left(\xi x\right)dx={\displaystyle\sum\limits_{j=0}^{\infty}}J(p^{-j})\left\{\text{ }{\displaystyle\int\limits_{p^{j}\mathbb{Z}_{p}}}\chi_{p}\left(\xi x\right)dx-{\displaystyle\int\limits_{p^{j+1}\mathbb{Z}_{p}}}\chi_{p}\left(\xi x\right)dx\right\}.$
Now, the announced formula follows from (9.1), by the following formula:
${\displaystyle\int\limits_{p^{j}\mathbb{Z}_{p}}}\chi_{p}\left(\xi x\right)dx=p^{-j}{\displaystyle\int\limits_{\mathbb{Z}_{p}}}\chi_{p}\left(p^{j}\xi y\right)dy=p^{-j}\Omega\left(p^{-j}\left|\xi\right|_{p}\right)\text{, where }x=p^{j}y,dx=p^{-j}dy.$
(ii) By changing variables as $y=ux$, $dy=dx$, and using that $\left|u^{-1}y\right|_{p}$ $\left|y\right|_{p}$, for any $u\in\mathbb{Z}_{p}^{\times}$,
$\widehat{J}(u\xi)={\displaystyle\int\limits_{\mathbb{Z}_{p}}}J(\left|x\right|_{p})\chi_{p}\left(\xi ux\right)dx={\displaystyle\int\limits_{\mathbb{Z}_{p}}}J(\left|y\right|_{p})\chi_{p}\left(\xi y\right)dy=\widehat{J}(\xi).$
(iii) If $\xi_{0}\in\mathbb{Z}_{p}$, then $\chi_{p}\left(\xi_{0}x\right)=1$ for any $x\in\mathbb{Z}_{p}$, consequently, $\widehat{J}(\xi)=\widehat{J}(\xi+\xi_{0})$ for any $\xi\in\mathbb{Z}_{p}$. Finally, for $\xi_{0}\in\mathbb{Z}_{p}$,
$\widehat{J}(\xi_{0})={\displaystyle\int\limits_{\mathbb{Z}_{p}}}J(\left|x\right|_{p})dx=\left\|J(\left|x\right|_{p})\right\|_{1}=\widehat{J}(0).$
Remark 9.2.
As a corollary, the function $F(\xi)=\exp\left(-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)\right)$, for $t>0$, is a locally constant function in $\mathbb{Q}_{p}$, satisfying $F(\xi)=F(\xi+\xi_{0})$, for $\xi\in\mathbb{Q}_{p}$, $\xi_{0}\in\mathbb{Z}_{p}$. Then, $F(\xi)$ defines a distribution from $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$, whose Fourier transform is a distribution supported in the unit ball,
$Z_{0}\left(x,t\right):=\left\{\begin{array}[c]{ll}\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}\right)\in\mathcal{D}^{\prime}\left(\mathbb{Z}_{p}\right),&\text{if }t>0.\\ &\\ \delta\left(x\right)&\text{if }t=0,\end{array}\right$.
where $\delta\left(x\right)$ is the Dirac distribution, see, e.g., [19, Section VII.3], [35, Theorem 4.9.3], [32, Theorem 1.161]. Notice that $\left|\widehat{J}(\left|\xi\right|_{p})\right|\leq 1$, for any $\xi\in\mathbb{Q}_{p}$.
Lemma 9.2.
With the above notation, the following formula holds:
(9.2) $Z_{0}\left(x,t\right)$ $=\Omega\left(\left|x\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{\infty}}\mathrm{e}^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}\left\{p^{j}\Omega\left(p^{j}\left|x\right|_{p}\right)-p^{\left(j-1\right)}\Omega\left(p^{\left(j-1\right)}\left|x\right|_{p}\right)\right\}$
$=\Omega\left(\left|x\right|_{p}\right)-\mathrm{e}^{-t\left(1-\widehat{J}\left(p\right)\right)}1_{\mathbb{Z}_{p}^{\times}}\left(x\right)+\Omega\left(p\left|x\right|_{p}\right){\displaystyle\sum\limits_{j=1}^{\infty}}p^{j}\mathrm{e}^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)},$
for $x\in\mathbb{Z}_{p}\smallsetminus\left\{0\right\}$, and $t>0$.
Remark 9.3.
We take $x\neq 0$, due to the fact, that for x=0\ x=0, the convergence of the series
$\left(1-p^{-1}\right){\displaystyle\sum\limits_{j=1}^{\infty}}p^{j}\mathrm{e}^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}$
is unknown, i.e., we require an extra hypothesis on $\widehat{J}$ to determine it. Also, for $x\neq 0$, the series in (9.2) is a finite sum due to the fact that $p^{j}\Omega\left(p^{j}\left|x\right|_{p}\right)-p^{\left(j-1\right)}\Omega\left(p^{\left(j-1\right)}\left|x\right|_{p}\right)=0$ for $j\geq 2+ord(x)$.
Proof.
By using that $\Omega\left(p^{-j}\left|x\right|_{p}\right)-\Omega\left(p^{-\left(j-1\right)}\left|x\right|_{p}\right)$ is the characteristic function of the set $S_{j}=p^{-j}\mathbb{Z}_{p}\smallsetminus p^{-\left(j-1\right)}\mathbb{Z}_{p}$, we have
$\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}$ $=\Omega\left(\left|\xi\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{\infty}}\mathrm{e}^{-t\left(1-\widehat{J}(p^{j})\right)}1_{S_{j}}\left(\xi\right)$
$=\Omega\left(\left|\xi\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{\infty}}\mathrm{e}^{-t\left(1-\widehat{J}(p^{j})\right)}\left\{\Omega\left(p^{-j}\left|x\right|_{p}\right)-\Omega\left(p^{-\left(j-1\right)}\left|x\right|_{p}\right)\right\}.$
Now,
$\Omega\left(\left|\xi\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{m}}\mathrm{e}^{-t\left(1-\widehat{J}(p^{j})\right)}\left\{\Omega\left(p^{-j}\left|x\right|_{p}\right)-\Omega\left(p^{-\left(j-1\right)}\left|x\right|_{p}\right)\right\}\rightarrow\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}$
as $m\rightarrow\infty$ in $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$, and since the Fourier transform is a topological and algebraic isomorphism on $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$,
$\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\Omega\left(\left|\xi\right|_{p}\right)\right)+{\displaystyle\sum\limits_{j=1}^{m}}\mathrm{e}^{-t\left(1-\widehat{J}(p^{j})\right)}\mathcal{F}_{\xi\rightarrow x}^{-1}\left\{\Omega\left(p^{-j}\left|x\right|_{p}\right)-\Omega\left(p^{-\left(j-1\right)}\left|x\right|_{p}\right)\right\}\\ \rightarrow\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}\right)$
as $m\rightarrow\infty$ in $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$. Now, the announced formula follows from
$\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\Omega\left(\left|\xi\right|_{p}\right)\right)+{\displaystyle\sum\limits_{j=1}^{m}}\mathrm{e}^{-t\left(1-\widehat{J}(p^{j})\right)}\mathcal{F}_{\xi\rightarrow x}^{-1}\left\{\Omega\left(p^{-j}\left|x\right|_{p}\right)-\Omega\left(p^{-\left(j-1\right)}\left|x\right|_{p}\right)\right\}$
$=\Omega\left(\left|x\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{m}}\mathrm{e}^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}\left\{p^{j}\Omega\left(p^{j}\left|x\right|_{p}\right)-p^{\left(j-1\right)}\Omega\left(p^{\left(j-1\right)}\left|x\right|_{p}\right)\right\},$
by using that
$\mathcal{F}_{\xi\rightarrow x}^{-1}\left\{\Omega\left(p^{k}\left|x\right|_{p}\right)=p^{-k}\Omega\left(p^{-k}\left|x\right|_{p}\right)\right\}.$
The following result is established by using the above technique.
Lemma 9.3.
Set
$\boldsymbol{J}\left(Z_{0}\left(x,t\right)\right):=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\left\{\widehat{J}(\left|\xi\right|_{p})-1\right\}\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}\right)\in\mathcal{D}^{\prime}\left(\mathbb{Z}_{p}\right),$
for $t>0$. Then
$\boldsymbol{J}\left(Z_{0}\left(x,t\right)\right)={\displaystyle\sum\limits_{j=1}^{\infty}}\mathrm{e}^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}\left(\widehat{J}\left(p^{j}\right)-1\right)\left\{p^{j}\Omega\left(p^{j}\left|x\right|_{p}\right)-p^{\left(j-1\right)}\Omega\left(p^{\left(j-1\right)}\left|x\right|_{p}\right)\right\},$
for $x\in\mathbb{Z}_{p}\smallsetminus\left\{0\right\}$, and $t>0$.
Lemma 9.4.
${\displaystyle\int\limits_{\mathbb{Z}_{p}}}Z_{0}(x,t)dx=1\text{, for }t>0\text{.}$
Proof.
By Lemma (9.1), for $t>0$,
$\left(Z_{0}(x,t),\Omega\left(\left|x\right|_{p}\right)\right)$ $=\left(\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}\right),\Omega\left(\left|x\right|_{p}\right)\right)=\left(\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)},\Omega\left(\left|\xi\right|_{p}\right)\right)$
$={\displaystyle\int\limits_{\mathbb{Z}_{p}}}\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}d\xi=\mathrm{e}^{-t\left(1-\widehat{J}(0)\right)}{\displaystyle\int\limits_{\mathbb{Z}_{p}}}d\xi=1.$
Lemma 9.5.
(i) For $x\in I+p^{l}\mathbb{Z}_{p}$,
$Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right)=1-Z_{0}\left(1,t\right)\left(1-p^{-1}\right),$
for $t>0$.
(ii) For $x\in K+p^{l}\mathbb{Z}_{p}$, $K\neq I$ as elements from $G_{l}$,
$Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right)=p^{-l}Z_{0}\left(\left|K-I\right|_{p},t\right),$
for $t>0$.
(iii) For $x\in\mathbb{Z}_{p}$, and $t>0$,
$Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right)$ $=1-Z_{0}\left(1,t\right)\left(1-p^{-1}\right)\Omega\left(p^{l}\left|x-I\right|_{p}\right)+$
${\displaystyle\sum\limits_{\begin{subarray}{c}K\in G_{l}\\ K\neq I\end{subarray}}}p^{-l}Z_{0}\left(\left|K-I\right|_{p},t\right)\Omega\left(p^{l}\left|x-K\right|_{p}\right).$
(iv) The space $\mathcal{D}_{l}(\mathbb{Z}_{p})$ is invariant under the operator $\varphi\left(x\right)\rightarrow Z_{0}(x,t)\ast\varphi\left(x\right)$.
Proof.
Set
$Z_{0}^{\left(m\right)}\left(x,t\right)=\Omega\left(\left|x\right|_{p}\right)+{\displaystyle\sum\limits_{j=1}^{m}}\mathrm{e}^{-t\left(1-\widehat{J}\left(p^{j}\right)\right)}\left\{p^{j}\Omega\left(p^{j}\left|x\right|_{p}\right)-p^{\left(j-1\right)}\Omega\left(p^{\left(j-1\right)}\left|x\right|_{p}\right)\right\}.$
Then, $Z_{0}^{\left(m\right)}\left(x,t\right)\rightarrow Z_{0}\left(x,t\right)$ as $m\rightarrow\infty$ in $\mathcal{D}^{\prime}\left(\mathbb{Q}_{p}\right)$. We now recall that the convolution of a distribution with a test function is a locally constant function (see, e.g. [32, Theorem 1.154]) given by
$Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right)=\left(Z_{0}\left(y,t\right),\Omega\left(p^{l}\left|x-I-y\right|_{p}\right)\right)$
$=\lim_{m\rightarrow\infty}\left(Z_{0}^{\left(m\right)}\left(y,t\right),\Omega\left(p^{l}\left|x-I-y\right|_{p}\right)\right)=\lim_{m\rightarrow\infty}\text{ }{\displaystyle\int\limits_{x-I+p^{l}\mathbb{Z}p}}\text{ }Z_{0}^{\left(m\right)}\left(y,t\right)dy.$
We first consider the case, $x\in I+p^{l}\mathbb{Z}_{p}$, i.e. $x-I+p^{l}\mathbb{Z}_{p}=p^{l}\mathbb{Z}_{p}$,
$Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right)=\lim_{m\rightarrow\infty}\text{ }{\displaystyle\int\limits_{p^{l}\mathbb{Z}p}}\text{ }Z_{0}^{\left(m\right)}\left(y,t\right)dy={\displaystyle\int\limits_{p^{l}\mathbb{Z}p}}\text{ }Z_{0}\left(y,t\right)dy.$
Now, we use that $Z_{0}\left(y,t\right)=$ $Z_{0}\left(\left|y\right|_{p},t\right)$, and Lemma 9.4,
$1={\displaystyle\int\limits_{\mathbb{Z}_{p}}}Z_{0}\left(\left|y\right|_{p},t\right)dy={\displaystyle\int\limits_{p^{l}\mathbb{Z}p}}\text{ }Z_{0}\left(\left|y\right|_{p},t\right)dy+{\displaystyle\int\limits_{\mathbb{Z}p^{\times}}}\text{ }Z_{0}\left(\left|y\right|_{p},t\right)dy$
for $t>0$, therefore
$Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right)={\displaystyle\int\limits_{p^{l}\mathbb{Z}p}}\text{ }Z_{0}\left(\left|y\right|_{p},t\right)dy=1-Z_{0}\left(1,t\right)\left(1-p^{-1}\right),$
for $t>0$.
We now consider the case, $x\in K+p^{l}\mathbb{Z}_{p}$, i.e., $K\neq I$ in $G_{l}=p^{l}\mathbb{Z}_{p}$, by using that $Z_{0}^{\left(m\right)}\left(y,t\right)=Z_{0}^{\left(m\right)}\left(\left|y\right|_{p},t\right)$,
$Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right)=\lim_{m\rightarrow\infty}\text{ }{\displaystyle\int\limits_{p^{l}\mathbb{Z}p}}\text{ }Z_{0}^{\left(m\right)}\left(y,t\right)dy=p^{-l}Z_{0}\left(\left|K-I\right|_{p},t\right).$
9.2. $p$-Adic heat equations and Markov processes in the unit ball
With $J$ as before, the mapping
$\begin{array}[c]{cccc}\boldsymbol{J}:&L^{\rho}\left(\mathbb{Z}_{p}\right)&\rightarrow&L^{\rho}\left(\mathbb{Z}_{p}\right)\\ &&&\\ &\varphi\left(x\right)&\rightarrow&J\left(\left|x\right|_{p}\right)\ast\varphi\left(x\right)-\varphi\left(x\right)\end{array}$
gives rise to a well-defined linear, bounded operator, for $\rho\in\left[0,\infty\right]$. Indeed,
$\left\|\boldsymbol{J}\varphi\left(x\right)\right\|_{\rho}=\left\|J\ast\varphi-\varphi\right\|_{\rho}\leq\left\|J\ast\varphi\right\|_{\rho}+\left\|\varphi\right\|_{\rho}\leq\left\|J\right\|_{1}\left\|\varphi\right\|_{\rho}+\left\|\varphi\right\|_{\rho}=2\left\|\varphi\right\|_{\rho}.$
Furthermore, $\boldsymbol{J}$ is a pseudo-differential operator,
$\boldsymbol{J}\varphi\left(x\right)=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\left(\widehat{J}(\left|\xi\right|_{p})-1\right)\mathcal{F}_{x\rightarrow\xi}\varphi\right)={\displaystyle\int\limits_{\mathbb{Q}_{p}}}\left(\widehat{J}(\left|\xi\right|_{p})-1\right)\widehat{\varphi}\left(\xi\right)\chi_{p}\left(-\xi x\right)dx,$
for $\varphi\in\mathcal{D}\left(\mathbb{Z}_{p}\right)$. Notice that $\boldsymbol{J}\varphi\in\mathcal{D}\left(\mathbb{Z}_{p}\right)$, indeed, $\widehat{J}(\left|\xi\right|_{p})-1$ is a locally constant function, then its product with $\widehat{\varphi}\left(\xi\right)$ is a Now, we use the fact that the space of the test function is invariant under the Fourier transform.
Proposition 9.1.
The function
$u\left(x,t\right)=Z_{0}\left(x,t\right)\ast u_{0}\left(x\right)\text{, }x\in\mathbb{Z}_{p},t\geq 0\text{,}$
is a solution of the Cauchy problem
(9.3) $\left\{\begin{array}[c]{l}u\left(\cdot,t\right)\in\mathcal{D}(\mathbb{Z}_{p})\text{ for }t\geq 0\text{, }u\left(x,\cdot\right)\in C^{1}\left[0,\infty\right)\\ \\ \frac{\partial}{\partial t}u\left(x,t\right)=\boldsymbol{J}u\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\geq 0\\ \\ u\left(x,0\right)=u_{0}\left(x\right)\in\mathcal{D}(\mathbb{Z}_{p}).\end{array}\right$.
Proof.
Since $u_{0}\left(x\right)\in\mathcal{D}(\mathbb{Z}_{p})=\cup_{l}\mathcal{D}_{l}(\mathbb{Z}_{p})$, we may assume that $u_{0}\left(x\right)\in\mathcal{D}_{l}(\mathbb{Z}_{p})$ for some positive integer $l$. on the other hand, $\frac{\partial}{\partial t}$, $\boldsymbol{J}$ are linear operators, and $\left\{\Omega\left(p^{l}\left|x-I\right|_{p}\right)\right\}_{I\in G_{l}}$ is a basis for $\mathcal{D}_{l}(\mathbb{Z}_{p})$, we may assume without loss of generality that $u_{0}\left(x\right)=\Omega\left(p^{l}\left|x-I\right|_{p}\right)$.
We now show that $u\left(x,t\right)=Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right)$, $x\in\mathbb{Z}_{p},t>0$, is a solution of the differential equation in (9.3). Indeed, by Lemmas 9.3, 9.5,
$\frac{\partial}{\partial t}u\left(x,t\right)=\frac{\partial}{\partial t}\left\{Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right)\right\}=\left(\frac{\partial}{\partial t}Z_{0}\left(x,t\right)\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right),$
in $\mathcal{D}^{\prime}\left(\mathbb{Z}_{p}\right)$. Now, by Lemmas 9.2 and 9.3,
$\frac{\partial}{\partial t}Z_{0}\left(x,t\right)=\boldsymbol{J}\left(Z_{0}\left(x,t\right)\right),$
in $\mathcal{D}^{\prime}\left(\mathbb{Z}_{p}\right)$, and by convoluting both sides with $\Omega\left(p^{l}\left|x-I\right|_{p}\right)$, we get
$\frac{\partial}{\partial t}u\left(x,t\right)=\boldsymbol{J}\left(Z_{0}\left(x,t\right)\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right).$
On the other hand,
$\boldsymbol{J}\left(Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right)\right)=\\ \mathcal{F}_{\xi\rightarrow x}^{-1}\left(\left\{\widehat{J}(\left|\xi\right|_{p})-1\right\}e^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}p^{-l}\chi_{p}\left(\xi I\right)\Omega\left(p^{-l}\left|x\right|_{p}\right)\right)\\ =\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\left\{\widehat{J}(\left|\xi\right|_{p})-1\right\}e^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right)\\ =\boldsymbol{J}\left(Z_{0}\left(x,t\right)\right)\ast\Omega\left(p^{l}\left|x-I\right|_{p}\right),$
in $\mathcal{D}^{\prime}\left(\mathbb{Z}_{p}\right)$, i.e., for $t>0$,
$\frac{\partial}{\partial t}u\left(x,t\right)=\boldsymbol{J}u\left(x,t\right)\text{ in }\mathcal{D}^{\prime}\left(\mathbb{Z}_{p}\right).$
The initial condition in (9.3) is satisfied because $Z_{0}\left(x,0\right)=\delta\left(x\right)$. Finally, $Z_{0}\left(x,t\right)\ast u_{0}\left(x\right)$ is a test function for every $t\geq 0$, cf. Lemma 9.5-(iii). ∎
We now review a result from ([40, Theorem 3.1]). We denote by $\mathcal{C}(\mathbb{Z}_{p})$ the $\mathbb{R}$-vector space of continuous functions on the unit ball. We recall that $\mathcal{D}\left(\mathbb{Z}_{p}\right)$ is dense in $\mathcal{C}(\mathbb{Z}_{p})$. We fix a time horizon $T\in\left(0,\infty\right]$, and consider the following initial value problem:
(9.4) $\left\{\begin{array}[c]{l}u\left(\cdot,t\right)\in\mathcal{C}(\mathbb{Z}_{p})\text{ for }t\geq 0\text{, }u\left(x,\cdot\right)\in\mathcal{C}^{1}\left(\left[0,T\right]\right)\text{ for }x\in\mathbb{Z}_{p}\\ \\ \frac{\partial}{\partial t}u\left(x,t\right)=J\left(\left|x\right|_{p}\right)\ast u\left(x,t\right)-u\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\in\left[0,T\right]\\ \\ u\left(x,0\right)=u_{0}\left(x\right)\in\mathcal{C}(\mathbb{Z}_{p}).\end{array}\right$.
The operator $\boldsymbol{J}\varphi\left(x\right)=J\left(\left|x\right|_{p}\right)\ast\varphi\left(x\right)-\varphi\left(x\right)$, for $\varphi\left(x\right)\in\mathcal{C}(\mathbb{Z}_{p})$, satisfies the maximum principle, i.e., if $\varphi\left(x_{0}\right)=\sup_{x\in\mathbb{Z}_{p}}\varphi\left(x\right)$, then
$\boldsymbol{J}\varphi\left(x_{0}\right)={\displaystyle\int\limits_{\mathbb{Z}_{p}}}J\left(\left|x_{0}-y\right|_{p}\right)\left\{\varphi\left(y\right)-\varphi\left(x_{0}\right)\right\}dy\leq 0\text{.}$
The condition $J\left(\left|x\right|_{p}\right)\leq 1$, $x\in\mathbb{Z}_{p}$, was used in ([40, Theorem 3.1]) to assure that operator $\boldsymbol{J}$ satisfies the maximum principle. Consequently, here we do not need this condition. By ([40, Theorem 3.1]), there exists a probability measure $p_{t}\left(x,\cdot\right)$, $t\in\left[0,T\right]$, with $T=T(u_{0})$, $x\in\mathbb{Z}_{p}$, on the Borel $\sigma$-algebra $\mathcal{B}$ of $\mathbb{Z}_{p}$, such that the Cauchy problem (9.4) has a unique solution of the form
$u(x,t)=\int\limits_{\mathbb{Z}_{p}}u_{0}(y)p_{t}\left(x,dy\right).$
In addition, $p_{t}\left(x,\cdot\right)$ is the transition function of a Markov process $\mathfrak{X}$ whose paths are right continuous and have no discontinuities other than jumps. We now determine the measure $p_{t}\left(x,\cdot\right)$.
By comparing the solutions of Cauchy problems (9.3) and (9.4), we conclude that
$\int\limits_{\mathbb{Z}_{p}}1_{B}(y)p_{t}\left(x,dy\right)=\int\limits_{\mathbb{Z}_{p}}1_{B}(y)Z_{0}\left(x-y,t\right)dy\text{, }$
for $t\in\left[0,T\right]$ and any open compact subset $B$ of $\mathbb{Z}_{p}$. Then $p_{t}\left(x,dy\right)$ exists for any $t\in\left[0,\infty\right]$, and since the open compact subset $B$ of $\mathbb{Z}_{p}$ generate the Borel $\sigma$-algebra $\mathcal{B}$ of $\mathbb{Z}_{p}$, then $p_{t}\left(x,dy\right)=Z_{0}\left(x-y,t\right)dy$ for $t\geq 0$. So we have the following result:
Theorem 9.1.
The Cauchy problem
(9.5) $\left\{\begin{array}[c]{l}u\left(\cdot,t\right)\in\mathcal{C}(\mathbb{Z}_{p})\text{ for }t\geq 0\text{, }u\left(x,\cdot\right)\in C^{1}\left[0,\infty\right)\\ \\ \frac{\partial}{\partial t}u\left(x,t\right)=J\left(\left|x\right|_{p}\right)\ast u\left(x,t\right)-u\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\geq 0\\ \\ u\left(x,0\right)=u_{0}\left(x\right)\in\mathcal{C}(\mathbb{Z}_{p}).\end{array}\right$.
has a unique solution of the form $u\left(x,t\right)=Z_{0}\left(x,t\right)\ast u_{0}\left(x\right)$. Furthermore,
(9.6) $p(t,x,B)=\begin{cases}\int_{B}Z_{0}(x-y,t)dy=1_{B}\left(x\right)\ast Z_{0}(x,t)&\text{for }t>0,\quad x\in\mathbb{Z}_{p},\quad B\in\mathcal{B}\\ 1_{B}(x)&\text{for }t=0,\end{cases}$
is the transition function of a Markov process $\mathfrak{X}$ whose paths are right continuous and have no discontinuities other than jumps. Furthermore,
${\displaystyle\int\limits_{\mathbb{Z}_{p}}}Z_{0}(x,t)dx=1\text{, for }t>0.$
Remark 9.4.
The value $p(t,x,B)$ gives the transition probability, per unit of time, that a particle/walker starting at the point $x\in\mathbb{Z}_{p}$ will be found in the set $B$ at the time $t$.
10. Appendix C: Continuous-time Markov chains
We take $\mathcal{K}_{r}$, $r\in\mathbb{J}$ as in (3.6), where $\mathbb{J}$ is a countable set, possibly infinite. The value
$p_{r,v}(t):={\displaystyle\int\limits_{\mathcal{K}_{v}}}p(t,x,\mathcal{K}_{r})\frac{dx}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}={\displaystyle\int\limits_{\mathcal{K}_{v}}}Z_{0}\left(x,t\right)\ast 1_{\mathcal{K}_{r}}\left(x\right)\frac{dx}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)},$
see (9.6), gives the transition probability, per unit of time, that a particle/walker starting at some position in $\mathcal{K}_{v}$ will be found in the set $\mathcal{K}_{r}$ at the time $t$. Notice that
$p_{r,v}(0)=\left\{\begin{array}[c]{ccc}1&\text{if}&r=v\\ &&\\ 0&\text{if}&r\neq v.\end{array}\right$.
Lemma 10.1.
With the above notation, for any $v\in\mathbb{J}$ , for $t>0$,
${\displaystyle\sum\limits_{r\in\mathbb{J}}}p_{r,v}(t)=1.$
Proof.
We fix a family $\left\{\mathbb{J}_{n}\right\}_{n\in\mathbb{N}}$ of subsets of $\mathbb{J}$ such that each $\mathbb{J}_{n}$ is a finite set, $\mathbb{J}_{n}\subset\mathbb{J}_{n+1}$ for every $n\in\mathbb{N}$, and $\cup_{n\in\mathbb{N}}\mathbb{J}_{n}=\mathbb{J}$. Notice that
$Z_{0}\left(x,t\right)\ast{\displaystyle\sum\limits_{r\in\mathbb{J}_{n}}}1_{\mathcal{K}_{r}}\left(x\right)$ $={\displaystyle\int\limits_{\mathbb{Z}_{p}}}Z_{0}\left(y,t\right)\left({\displaystyle\sum\limits_{r\in\mathbb{J}_{n}}}1_{\mathcal{K}_{r}}\left(x-y\right)\right)dy\leq\Omega\left(\left|x\right|_{p}\right){\displaystyle\int\limits_{\mathbb{Z}_{p}}}Z_{0}\left(y,t\right)dy$
$\leq\mathrm{e}^{-t\left(1-\widehat{J}(0)\right)}\Omega\left(\left|x\right|_{p}\right)=\Omega\left(\left|x\right|_{p}\right)\text{, }$
since $\widehat{J}(0)=1$, for any $t>0$, cf. Lemma 9.4. By the dominated convergence theorem,
$\sum\limits_{r\in\mathbb{J}}p_{r,v}(t)=\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\lim_{n\rightarrow\infty}\sum\limits_{r\in\mathbb{J}_{n}}\text{ }\int\limits_{\mathcal{K}_{v}}Z_{0}\left(x,t\right)\ast 1_{\mathcal{K}_{r}}\left(x\right)dx\\ =\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\lim_{n\rightarrow\infty}\int\limits_{\mathcal{K}_{v}}Z_{0}\left(x,t\right)\ast\sum\limits_{r\in\mathbb{J}_{n}}1_{\mathcal{K}_{r}}\left(x\right)dx\\ =\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\int\limits_{\mathcal{K}_{v}}Z_{0}\left(x,t\right)\ast\lim_{n\rightarrow\infty}\sum\limits_{r\in\mathbb{J}_{n}}1_{\mathcal{K}_{r}}\left(x\right)dx=\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\int\limits_{\mathcal{K}_{v}}Z_{0}\left(x,t\right)\ast\Omega\left(\left|x\right|_{p}\right)dx\\ =\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\int\limits_{\mathcal{K}_{v}}\left\{\int\limits_{\mathbb{Z}_{p}}Z_{0}\left(x-y,t\right)dy\right\}dx=\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\int\limits_{\mathcal{K}_{v}}\left\{\int\limits_{\mathbb{Z}_{p}}Z_{0}\left(u,t\right)du\right\}dx\\ =\frac{\mathrm{e}^{-t\left(1-\widehat{J}(0)\right)}\int_{\mathcal{K}_{v}}dx}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}=1.$
Theorem 10.1.
$\left[p_{r,v}(t)\right]_{r,v\in\mathbb{J}}$ is the transition matrix of a CTMC. If $\widehat{J}(\left|u\right|_{p})<1$, for $u\in\mathbb{Q}_{p}\smallsetminus\mathbb{Z}_{p}$, then
$\lim_{t\rightarrow\infty}\left|p_{r,v}(t)-\mu_{\text{Haar}}(\mathcal{K}_{r})\right|=0\text{,}$
for $t>0$. Consequently, the CTMC admits a stationary distribution $\left[p_{r}^{\text{sta}}\right]_{r\in\mathbb{J}}$, where
$p_{r}^{\text{sta}}=\lim_{t\rightarrow\infty}p_{r,v}(t)=\mu_{\text{Haar}}(\mathcal{K}_{r}).$
Proof.
(i) By using that $\mathbb{Z}_{p}$ is an additive topological group, for a fixed $x\in\mathbb{Z}_{p}$, the mapping $y\rightarrow z=x-y$ is a topological and algebraic isomorphism preserving the Haar measure, then
$\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}{\displaystyle\int\limits_{\mathcal{K}_{v}}}\Omega\left(\left|x\right|_{p}\right)\ast 1_{\mathcal{K}_{r}}\left(x\right)dx=\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}{\displaystyle\int\limits_{\mathcal{K}_{v}}}{\displaystyle\int\limits_{\mathbb{Z}_{p}}}1_{\mathcal{K}_{r}}\left(x-y\right)dy$
(10.1) $=\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}{\displaystyle\int\limits_{\mathcal{K}_{v}}}{\displaystyle\int\limits_{\mathbb{Z}_{p}}}1_{\mathcal{K}_{r}}\left(z\right)dz=\frac{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)\mu_{\text{Haar}}\left(\mathcal{K}_{r}\right)}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}=\mu_{\text{Haar}}(\mathcal{K}_{r}).$
Now, for $t>0$, we set
$F(x;t):=1_{\mathcal{K}_{v}}\left(x\right)\left\{\left\{Z_{0}\left(x,t\right)-\Omega\left(\left|x\right|_{p}\right)\right\}\ast 1_{\mathcal{K}_{r}}\left(x\right)\right\}.$
Since the convolution of a distribution ($Z_{0}\left(x,t\right)-\Omega\left(\left|x\right|_{p}\right)$) and a test function ($1_{\mathcal{K}_{r}}\left(x\right)$) is a locally function; and that the product of this function times the test function $1_{\mathcal{K}_{v}}\left(x\right)$ is a test function, then $F(x)$ is a test function in $x$, for $t>0$. Therefore,
${\displaystyle\int\limits_{\mathbb{Q}_{p}}}F(x,t)dx=\widehat{F}(0,t).$
On the other hand, by using (10.1),
$\left|p_{r,v}(t)-\mu_{\text{Haar}}(\mathcal{K}_{r})\right|$ $=\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\left|\text{ }{\displaystyle\int\limits_{\mathcal{K}_{v}}}\left\{Z_{0}\left(x,t\right)-\Omega\left(\left|x\right|_{p}\right)\right\}\ast 1_{\mathcal{K}_{r}}\left(x\right)dx\right|$
$=\frac{1}{\mu_{\text{Haar}}\left(\mathcal{K}_{v}\right)}\left|\text{ }{\displaystyle\int\limits_{\mathbb{Q}_{p}}}F(x,t)dx\right|=\left|\widehat{F}(0,t)\right|.$
Then, we have to show that $\lim_{t\rightarrow\infty}\widehat{F}(0,t)=0.$
By using $1=\widehat{J}(0)$ and $\widehat{J}(\left|\xi\right|_{p})=\widehat{J}(0)$ for $\left|\xi\right|_{p}\leq 1$,
$\widehat{F}(\xi,t)$ $=\widehat{1}_{\mathcal{K}_{v}}\left(\xi\right)\ast\left\{\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}-\Omega\left(\left|\xi\right|_{p}\right)\right\}\widehat{1}_{\mathcal{K}_{r}}\left(\xi\right)$
$=\widehat{1}_{\mathcal{K}_{v}}\left(\xi\right)\ast\left\{\begin{array}[c]{lll}0&\text{if }&\left|\xi\right|_{p}\leq 1\\ &&\\ \widehat{1}_{\mathcal{K}_{r}}\left(\xi\right)\mathrm{e}^{-t\left(1-\widehat{J}(\left|\xi\right|_{p})\right)}&\text{if }&\left|\xi\right|_{p}\geq p\end{array}\right$.
$={\displaystyle\int\limits_{\mathbb{Q}_{p}\smallsetminus\mathbb{Z}_{p}}}\widehat{1}_{\mathcal{K}_{v}}\left(\xi-u\right)\mathrm{e}^{-t\left(1-\widehat{J}(\left|u\right|_{p})\right)}\widehat{1}_{\mathcal{K}_{r}}\left(u\right)du.$
Now, by the hypothesis $1-\widehat{J}(\left|u\right|_{p})>0$ for $u\in\mathbb{Q}_{p}\smallsetminus\mathbb{Z}_{p}$, $\mathrm{e}^{-t\left(1-\widehat{J}(\left|u\right|_{p})\right)}\leq 1$ for $t>0$, and the fact that $\widehat{1}_{\mathcal{K}_{v}}\left(-u\right)\widehat{1}_{\mathcal{K}_{r}}\left(u\right)$ is a test function, we can apply the dominated convergence theorem,
$\lim_{t\rightarrow\infty}\widehat{F}(0,t)$ $=\lim_{t\rightarrow\infty}{\displaystyle\int\limits_{\mathbb{Q}_{p}\smallsetminus\mathbb{Z}_{p}}}\widehat{1}_{\mathcal{K}_{v}}\left(-u\right)\mathrm{e}^{-t\left(1-\widehat{J}(\left|u\right|_{p})\right)}\widehat{1}_{\mathcal{K}_{r}}\left(u\right)du$
$={\displaystyle\int\limits_{\mathbb{Q}_{p}\smallsetminus\mathbb{Z}_{p}}}\widehat{1}_{\mathcal{K}_{v}}\left(-u\right)\left\{\lim_{t\rightarrow\infty}\mathrm{e}^{-t\left(1-\widehat{J}(\left|u\right|_{p})\right)}\right\}\widehat{1}_{\mathcal{K}_{r}}\left(u\right)du=0.$
11. Appendix D: $p$-adic Schrödinger equations on the unit ball and quantum Markov chains
By performing a Wick rotation $t\rightarrow it$ in (9.5), and taking $\Psi\left(x,t\right)=u\left(x,it\right)$, we obtain the Schrödinger equations in the unit ball,
(11.1) $\left\{\begin{array}[c]{l}\mathrm{i}\frac{\partial}{\partial t}\Psi\left(x,t\right)=-J\left(\left|x\right|_{p}\right)\ast\Psi\left(x,t\right)+\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\geq 0\\ \\ \Psi\left(x,0\right)=\psi_{0}\left(x\right)\in\mathcal{D}(\mathbb{Z}_{p}).\end{array}\right$.
By using Fourier transforms, we obtain that the solution of (11.1) is given by
(11.2) $\Psi\left(x,t\right)=\mathcal{F}_{\xi\rightarrow x}^{-1}(\mathrm{e}^{-\mathrm{i}\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t})\ast\psi_{0}\left(x\right)\text{ in }\mathcal{D}^{\prime}(\mathbb{Q}_{p})\text{,}$
where $\mathcal{F}_{\xi\rightarrow x}$ denotes the Fourier transform in $\mathcal{D}^{\prime}(\mathbb{Q}_{p})$. Notice that $\Psi\left(x,t\right)$ is a locally constant function in $x$, for any $t\geq 0$.
Since
$\mathrm{e}^{-i\left(1-\widehat{J}\left(\left|\xi+\xi_{0}\right|_{p}\right)\right)t}=\mathrm{e}^{-i\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t}$
for any $\xi_{0}\in\mathbb{Z}_{p}$, $\xi\in\mathbb{Q}_{p}$, $\mathcal{F}_{\xi\rightarrow x}^{-1}(\mathrm{e}^{-\mathrm{i}\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t})$ is a distribution with support in $\mathbb{Z}_{p}$, see, e.g., [19, Section VII.3], [35, Theorem 4.9.3], [32, Theorem 1.161]. The explicit calculation of $\mathcal{F}_{\xi\rightarrow x}^{-1}(\mathrm{e}^{-\mathrm{i}\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t})$ can be carried out as in the case of the heat kernel, and one obtains that
$\mathcal{F}_{\xi\rightarrow x}^{-1}(\mathrm{e}^{-\mathrm{i}\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t}))=Z_{0}(x,\mathrm{i}t)$
for $x\in\mathbb{Z}_{p}\smallsetminus\left\{0\right\}$, $t>0$.
The formula (11.2) is useful; however, we need to show that $\left\|\Psi\left(\cdot,t\right)\right\|_{2}=1$, for any $\left\|\psi_{0}\right\|_{2}=1$, and any $t\geq 0$. To achieve this result, we need a different approach.
Lemma 11.1.
(i) For $k\in\left\{1,\ldots,p-1\right\}$, $r\leq 0$, and $b\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$, $bp^{-r}\in\mathbb{Z}_{p}$,
$\boldsymbol{J}\Psi_{rbk}\left(x\right)=(\widehat{J}\left(p^{1-r}\right)-1)\Psi_{rbk}\left(x\right).$
$\boldsymbol{J}\Omega\left(\left|x\right|_{p}\right)=0\text{.}$
Proof.
(i) We recall that $\boldsymbol{J}$ is a pseudo-differential operator,
$\boldsymbol{J}\Psi_{rbk}\left(x\right)=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\left(\widehat{J}\left(\left|\xi\right|_{p}\right)-1\right)\mathcal{F}_{x\rightarrow\xi}\Psi_{rbk}\left(x\right)\right).$
Now
$\widehat{\Psi}_{rbk}\left(\xi\right)=p^{\frac{r}{2}}\chi_{p}\left(p^{-r}b\xi\right)\Omega\left(\left|p^{-r}\xi+p^{-1}k\right|_{p}\right),$
where $\Omega\left(\left|p^{-r}\xi+p^{-1}k\right|_{p}\right)=1$ if and only if
$\xi\in-p^{r-1}k+p^{r}\mathbb{Z}_{p}=p^{r-1}k+p^{r}\mathbb{Z}_{p}.$
Then, by the ultrametric property of $\left|\cdot\right|_{p}$,
$\left|\xi\right|_{p}=\left|p^{r-1}k+p^{r}m\right|_{p}=\left|p^{r-1}k\right|_{p}=p^{1-r},$
for any $m\in\mathbb{Z}_{p}$. Then,
$\mathcal{F}_{x\rightarrow\xi}\left\{\boldsymbol{J}\left(\Psi_{rbk}\left(x\right)\right)\right\}=(\widehat{J}\left(p^{1-r}\right)-1)\widehat{\Psi}_{rbk}\left(\xi\right).$
(ii) The second part follows from
$\boldsymbol{J}\Omega\left(\left|x\right|_{p}\right)=J\left(\left|x\right|_{p}\right)\ast\Omega\left(\left|x\right|_{p}\right)-\Omega\left(\left|x\right|_{p}\right)$
$=\Omega\left(\left|x\right|_{p}\right)\int\limits_{\mathbb{Z}_{p}}J\left(\left|x-y\right|_{p}\right)\Omega\left(\left|y\right|_{p}\right)dy-\Omega\left(\left|x\right|_{p}\right)$
$=\Omega\left(\left|x\right|_{p}\right)\int\limits_{\mathbb{Z}_{p}}J\left(\left|y\right|_{p}\right)dy-\Omega\left(\left|x\right|_{p}\right)$
$=\Omega\left(\left|x\right|_{p}\right)\int\limits_{\mathbb{Z}_{p}}J\left(\left|y\right|_{p}\right)dy-\Omega\left(\left|x\right|_{p}\right)=\Omega\left(\left|x\right|_{p}\right)\left(\left\|J\right\|_{L^{1}\left(\mathbb{Z}_{p}\right)}-1\right).$
Theorem 11.1.
The initial value problem
[…]
has a unique solution of the form
(11.4) $\Psi(x,t)=A_{0}\Omega\left(\left|x\right|_{p}\right)+{\displaystyle\sum\limits_{k\in\left\{1,\ldots,p-1\right\}}}\text{ \ }{\displaystyle\sum\limits_{\begin{subarray}{c}b\in\mathbb{Q}_{p}/\mathbb{Z}_{p}\\ bp^{-r}\in\mathbb{Z}_{p}\end{subarray}}}\text{ \ }{\displaystyle\sum\limits_{r\leq 0}}\text{ }A_{{}_{rbk}}\mathrm{e}^{-\mathrm{i}\left(J\left(p^{1-r}\right)-1\right)t}\Psi_{rbk}\left(x\right){,}$
where the $A_{0}$, $A_{kr}\in\mathbb{C}$.
Proof.
The expansion (11.4) follows from the fact that $\left\{\Omega\left(\left|x\right|_{p}\right)\right\}\cup\left\{\Psi_{rbk}\left(x\right)\right\}$ is an orthonormal basis of $L^{2}\left(\mathbb{Z}_{p}\right)$, by using the separation of variables technique. Now,
$\left\|\Psi(\cdot,t)\right\|_{2}^{2}=\left|A_{0}\right|^{2}+{\displaystyle\sum\limits_{k\in\left\{1,\ldots,p-1\right\}}}\text{ \ }{\displaystyle\sum\limits_{\begin{subarray}{c}b\in\mathbb{Q}_{p}/\mathbb{Z}_{p}\\ bp^{-r}\in\mathbb{Z}_{p}\end{subarray}}}\text{ \ }{\displaystyle\sum\limits_{r\leq 0}}\text{ }\left|A_{{}_{rbk}}\right|^{2}=\left\|\psi_{0}\right\|_{2}^{2}=1.$
11.1. CTQMCs
We now use Theorem 3.1 to construct a family of CTQMCs. We use all the notation introduced in Subsections 3.3, 3.4,
Proposition 11.1.
For $l\geq 1$, set $\mathbb{J}=G_{l}$ and
$\pi_{r,v}\left(t\right)=\left|\left\langle p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right),Z_{0}(x,\mathrm{i}t)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\right\rangle\right|^{2},$
for $r,v\in G_{l}$. Then, $\left[\pi_{r,v}\left(t\right)\right]_{r,v\in G_{l}}$ is the transition matrix of a CTQMC.
Proof.
We use all the notation introduced in Subsection $3.$ Set
$e_{v}=p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right),$
for $v\in G_{l}$. Then $\left\{p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\right\}_{v\in G_{l}}$ is an orthonormal basis of the finite-dimensional Hilbert space $\mathcal{H}_{G_{l}}=\mathcal{D}_{l}(\mathbb{Z}_{p})$ Furthermore,
$\mathrm{e}^{\mathrm{i}t\boldsymbol{J}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)=Z_{0}(x,\mathrm{i}t)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\in\mathcal{D}_{l}(\mathbb{Z}_{p}),$
cf. Lemma 9.5-(iii). This fact implies that $\mathrm{e}^{\mathrm{i}t\boldsymbol{J}}\mathcal{D}_{l}(\mathbb{Z}_{p})\subset\mathcal{D}_{l}(\mathbb{Z}_{p})$. Now, the announced result follows from Theorem 3.1. ∎
Remark 11.1.
Using the wavefunctions $\Psi(x,t)$ and applying the techniques introduced in Subsection 3.4, we can construct a very large class of CTQMCs. However, the explicit computation of the $p$-adic wavefunctions, even for simple models, it is very involved, see [39]. Then, for practical purposes, we need good discretizations (in the sense of numerical analysis) of the wavefunctions. Our next step is to discuss the connection between this work and [26].
For a positive integer $l$, we set
$\alpha_{r,v}\left(t\right):=\left\langle p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right),Z_{0}(x,\mathrm{i}t)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\right\rangle,$
for $r,v\in G_{l}$, and $t\geq 0$. The $\alpha_{r,v}\left(t\right)$ are the (complex) transition amplitudes. Furthermore, $\pi_{r,v}\left(t\right)=\left|\alpha_{r,v}\left(t\right)\right|^{2}$.
Proposition 11.2.
With the above notation,
(11.5) $\alpha_{r,v}\left(t\right)=p^{-l}\left\{\mathrm{e}^{\mathrm{i}t}{\displaystyle\int\limits_{\left|\xi\right|_{p}\leq p^{l}}}\chi_{p}\left(\left(r-v\right)\xi\right)e^{-it\widehat{J}(\left|\xi\right|_{p})}d\xi\right\},$
for $r,v\in G_{l}$, and $t\geq 0$.
Proof.
By Theorem 11.1, in Appendix D,
$Z_{0}(x,\mathrm{i}t)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\in L^{2}\left(\mathbb{Z}_{p}\right)\subset L^{2}\left(\mathbb{Q}_{p}\right).$
Using the fact that the Fourier transform preserves the inner product in $L^{2}\left(\mathbb{Q}_{p}\right)$, we have
$\left\langle p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right),Z_{0}(x,\mathrm{i}t)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\right\rangle$
$=p^{-l}\left\langle\Omega\left(p^{-l}\left|\xi\right|_{p}\right)\chi_{p}\left(r\xi\right),\mathrm{e}^{\mathrm{i}t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\Omega\left(p^{-l}\left|\xi\right|_{p}\right)\chi_{p}\left(v\xi\right)\right\rangle$
$=p^{-l}{\displaystyle\int\limits_{\mathbb{Q}_{p}}}\Omega\left(p^{-l}\left|\xi\right|_{p}\right)\chi_{p}\left(r\xi\right)\mathrm{e}^{-\mathrm{i}t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\Omega\left(p^{-l}\left|\xi\right|_{p}\right)\chi_{p}\left(-v\xi\right)d\xi$
$=p^{-l}{\displaystyle\int\limits_{\left|\xi\right|_{p}\leq p^{l}}}\chi_{p}\left(\left(r-v\right)\xi\right)\mathrm{e}^{-\mathrm{i}t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}d\xi.$
We now summarize our results in the following theorem: ∎
Theorem 11.2.
With the notation introduced in Propositions 11.1, 11.2.
$\left[\pi_{r,v}\left(t\right)\right]_{r,v\in G_{l}}=\left[p^{-2l}\left|\text{ }{\displaystyle\int\limits_{\left|\xi\right|_{p}\leq p^{l}}}\chi_{p}\left(\left(r-v\right)\xi\right)\mathrm{e}^{-\mathrm{i}t\widehat{J}(\left|\xi\right|_{p})}d\xi\right|^{2}\right]_{r,v\in G_{l}}$
is the transition matrix of a CTQMC.
In particular, the $\lim_{t\rightarrow\infty}\pi_{r,v}\left(t\right)$ does not exist, and the announced CTQMCs do not admit a stationary distribution, nor do they have absorbing states.
12. Appendix E: Discretization of $p$-adic heat and Schrödinger equations
12.1. Some technical results
Lemma 12.1.
We use $J$ and $G_{l}=\mathbb{Z}_{p}/p^{l}\mathbb{Z}_{p}$ as before. Then
$\boldsymbol{J}\Omega\left(p^{l}\left|x-K\right|_{p}\right)$ $=J\left(x\right)\ast\Omega\left(p^{l}\left|x-K\right|_{p}\right)=p^{-l}{\displaystyle\sum\limits_{I\in G_{l}\smallsetminus\left\{0\right\}}}J\left(\left|I-K\right|_{p}\right)\Omega\left(p^{l}\left|x-I\right|_{p}\right)+$
$\left(\text{ }{\displaystyle\int\limits_{p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy\right)\Omega\left(p^{l}\left|x-K\right|_{p}\right),$
for $K\in G_{l}$.
Remark 12.1.
Notice that
${\displaystyle\int\limits_{\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy={\displaystyle\sum\limits_{I\in G_{l}}}\text{ \ }{\displaystyle\int\limits_{I+p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy={\displaystyle\sum\limits_{\begin{subarray}{c}I\in G_{l}\\ I\neq 0\end{subarray}}}p^{-l}J\left(\left|I\right|_{p}\right)+{\displaystyle\int\limits_{p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy,$
i.e.,
${\displaystyle\int\limits_{p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy={\displaystyle\int\limits_{\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy-{\displaystyle\sum\limits_{\begin{subarray}{c}I\in G_{l}\\ I\neq 0\end{subarray}}}p^{-l}J\left(\left|I\right|_{p}\right).$
Proof.
By using that
$\Omega\left(\left|x\right|_{p}\right)={\displaystyle\sum\limits_{I\in G_{l}}}\Omega\left(p^{l}\left|x-I\right|_{p}\right),$
and $J\left(\left|x\right|_{p}\right)\Omega\left(p^{l}\left|x-I\right|_{p}\right)=J\left(\left|I\right|_{p}\right)$, for $I\in G_{l}\smallsetminus\left\{0\right\}$, we have
$J\left(\left|x\right|_{p}\right)=\Omega\left(\left|x\right|_{p}\right)J\left(\left|x\right|_{p}\right)={\displaystyle\sum\limits_{I\in G_{l}\smallsetminus\left\{0\right\}}}J\left(\left|I\right|_{p}\right)\Omega\left(p^{l}\left|x-I\right|_{p}\right)+J\left(\left|x\right|_{p}\right)\Omega\left(p^{l}\left|x\right|_{p}\right).$
From this formula, by using the identity,
$\Omega\left(p^{l}\left|x-I\right|_{p}\right)\ast\Omega\left(p^{l}\left|x-K\right|_{p}\right)=p^{-l}\Omega\left(p^{l}\left|x-\left(I+K\right)\right|_{p}\right),$
we obtain
$J\left(\left|x\right|_{p}\right)\ast\Omega\left(p^{l}\left|x-K\right|_{p}\right)$ $=p^{-l}{\displaystyle\sum\limits_{I\in G_{l}\smallsetminus\left\{0\right\}}}J\left(\left|I\right|_{p}\right)\Omega\left(p^{l}\left|x-\left(I+K\right)\right|_{p}\right)+$
$\left\{J\left(\left|x\right|_{p}\right)\Omega\left(p^{l}\left|x\right|_{p}\right)\right\}\ast\Omega\left(p^{l}\left|x-K\right|_{p}\right).$
Set
$L\left(x\right):=\left\{J\left(\left|x\right|_{p}\right)\Omega\left(p^{l}\left|x\right|_{p}\right)\right\}\ast\Omega\left(p^{l}\left|x-K\right|_{p}\right)$
$={\displaystyle\int\limits_{\mathbb{Q}_{p}}}J\left(\left|y\right|_{p}\right)\Omega\left(p^{l}\left|y\right|_{p}\right)\Omega\left(p^{l}\left|x-K-y\right|_{p}\right)dy$
$={\displaystyle\int\limits_{x-K+p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)\Omega\left(p^{l}\left|y\right|_{p}\right)dy.$
The calculation of the last integral involves two cases.
Case 1: $x\in K+p^{l}\mathbb{Z}_{p}$
In this case $x-K\in p^{l}\mathbb{Z}_{p}$, and
$L\left(x\right)={\displaystyle\int\limits_{p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)\Omega\left(p^{l}\left|y\right|_{p}\right)dy=\left(\text{ }{\displaystyle\int\limits_{p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy\right)\Omega\left(p^{l}\left|x-K\right|_{p}\right).$
Case 2: $x\notin K+p^{l}\mathbb{Z}_{p}$
In this case, $\left(x-K+p^{l}\mathbb{Z}_{p}\right)\cap p^{l}\mathbb{Z}_{p}=\varnothing$, and
$L\left(x\right)={\displaystyle\int\limits_{x-K+p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)\Omega\left(p^{l}\left|y\right|_{p}\right)dy={\displaystyle\int\limits_{\left(x-K+p^{l}\mathbb{Z}_{p}\right)\cap p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy=0.$
In conclusion,
$J\left(\left|x\right|_{p}\right)\ast\Omega\left(p^{l}\left|x-K\right|_{p}\right)$ $=p^{-l}{\displaystyle\sum\limits_{I\in G_{l}\smallsetminus\left\{0\right\}}}J\left(\left|I\right|_{p}\right)\Omega\left(p^{l}\left|x-\left(I+K\right)\right|_{p}\right)+$
$\left({\displaystyle\int\limits_{p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy\right)\Omega\left(p^{l}\left|x-K\right|_{p}\right),$
and since $G_{l}$ is an additive group,
$J\left(x\right)\ast\Omega\left(p^{l}\left|x-K\right|_{p}\right)$ $=p^{-l}{\displaystyle\sum\limits_{I\in G_{l}\smallsetminus\left\{0\right\}}}J\left(\left|I-K\right|_{p}\right)\Omega\left(p^{l}\left|x-I\right|_{p}\right)+$
$\left({\displaystyle\int\limits_{p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy\right)\Omega\left(p^{l}\left|x-K\right|_{p}\right).$
Lemma 12.2.
The mapping $\boldsymbol{J}:\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)\rightarrow\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)$ is a well-defined operator. By identifying $\varphi\in\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)$ with column vector $\left[\varphi\left(I\right)\right]_{I\in G_{l}}$, the restriction of the operator $\boldsymbol{J}$ to $\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)$ is given by the matrix $\boldsymbol{J}^{\left(l\right)}=\left[J_{I,K}^{(l)}\right]_{I,K\in G_{l}}$, where
(12.1) $J_{I,K}^{(l)}=\left\{\begin{array}[c]{lll}p^{-l}J\left(\left|I-K\right|_{p}\right)&\text{if}&I\neq K\\ &&\\ \left(\text{ }{\displaystyle\int\limits_{p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy\right)-1&\text{if}&I=K.\end{array}\right$.
Remark 12.2.
We now consider $\boldsymbol{J}:L^{1}\left(\mathbb{Z}_{p}\right)\rightarrow L^{1}\left(\mathbb{Z}_{p}\right)$, then $\boldsymbol{J}^{\left(l\right)}=\boldsymbol{J}\mid_{\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)}$. For each $l\in\mathbb{N}$, there exists a linear bounded operator
$\boldsymbol{P}_{l}:L^{1}\left(\mathbb{Z}_{p}\right)\rightarrow\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right),$
satisfying $\boldsymbol{P}_{l}=\boldsymbol{P}_{l}^{2}$. For $f\in L^{1}\left(\mathbb{Z}_{p}\right)$, $\varphi_{l}=\boldsymbol{P}_{l}f$ is a good approximation of $f$ in $\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)$, i.e.,
$\left\|\varphi_{l}-f\right\|_{1}\rightarrow 0\text{ as }l\rightarrow\infty.$
Then $\boldsymbol{\boldsymbol{\boldsymbol{P}_{l}}J=J}^{\left(l\right)}$, and
$\left\|\boldsymbol{\boldsymbol{P}_{l}J}f-\boldsymbol{J}f\right\|_{1}=\left\|\boldsymbol{J}\varphi_{l}-\boldsymbol{J}f\right\|_{1}\leq\left\|\boldsymbol{J}\right\|\left\|\varphi_{l}-f\right\|_{1}\rightarrow 0\text{ as }l\rightarrow\infty.$
Which means, that $\boldsymbol{J}^{\left(l\right)}\rightarrow\boldsymbol{J}$, as $l\rightarrow\infty$, in the operator norm. See [26] for an in-depth discussion.
Remark 12.3.
Let $(\mathcal{G},V,E)$ be a simple graph; then, it does not contain any loops or multiple edges between the same pair of vertices, and all the edges are undirected. We take $V=G_{l}^{0}\subset G_{l}$ for some $l\in\mathbb{N}$ fixed. This means that we identify each vertex of $\mathcal{G}$ with a $p$-adic integer from $G_{l}$. Let $A=\left[A_{I,K}\right]_{I,K\in G_{l}^{0}}$ be the adjacency matrix of $\mathcal{G}$. Notice that $A_{I,I}=0$ since $\mathcal{G}$ has not loops, and $A_{I,K}=A_{K,I}$, for $I,K\in G_{l}^{0}$. For convenience’s sake, we assume that $A_{I,K}=0$ if $I\notin G_{l}^{0}$ or $K\notin G_{l}^{0}$. Now, we set
$J_{I,K}^{(l)}=p^{-l}A_{I,K}\text{ if }I\neq K\text{, with }I,K\in G_{l}^{0},$
and
$J_{I,K}^{(l)}=0\text{ if }I\notin G_{l}^{0}\text{ or }K\notin G_{l}^{0}.$
We assume that $1=\int_{\mathbb{Z}_{p}}J\left(\left|y\right|_{p}\right)dy$, then
(12.2) $1={\displaystyle\int\limits_{p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy+{\displaystyle\sum\limits_{\begin{subarray}{c}I\in G_{l}\\ I\neq 0\end{subarray}}}\text{ \ }{\displaystyle\int\limits_{I+p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy={\displaystyle\int\limits_{p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy+p^{-l}{\displaystyle\sum\limits_{\begin{subarray}{c}I\in G_{l}\\ I\neq 0\end{subarray}}}\text{ \ }J\left(\left|I\right|_{p}\right).$
Now, for any $K\in G_{l}$, the mapping
$\begin{array}[c]{ccc}G_{l}&\rightarrow&G_{l}\\ I&\rightarrow&I-K\end{array}$
is a group isomorphism. Then
(12.3) $p^{-l}{\displaystyle\sum\limits_{\begin{subarray}{c}I\in G_{l}\\ I\neq 0\end{subarray}}}\text{ \ }J\left(\left|I\right|_{p}\right)=p^{-l}{\displaystyle\sum\limits_{\begin{subarray}{c}I\in G_{l}\\ I\neq K\end{subarray}}}\text{ \ }J\left(\left|I-K\right|_{p}\right).$
Now from (12.2)-(12.3), taking $p^{-l}J\left(\left|I-K\right|_{p}\right)=p^{-l}A_{I,K}$, and using that
$val(K)=\sum_{\begin{subarray}{c}I\in G_{l}^{0}\\ I\neq K\end{subarray}}A_{I,K},$
where $val(K)$ denotes the valence of the vertex $K$, we have
${\displaystyle\int\limits_{p^{l}\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy$ $=1-p^{-l}{\displaystyle\sum\limits_{\begin{subarray}{c}I\in G_{l}\\ I\neq K\end{subarray}}}\text{ \ }J\left(\left|I-K\right|_{p}\right)=1-p^{-l}{\displaystyle\sum\limits_{\begin{subarray}{c}I\in G_{l}^{0}\\ I\neq K\end{subarray}}}\text{ }A_{I,K}$
$=1-p^{-l}val(K).$
In conclusion,
(12.4) $J_{I,K}^{(l)}=\left\{\begin{array}[c]{lll}p^{-l}A_{I,K}&\text{if}&I\neq K\\ &&\\ -p^{-l}val(K)&\text{if}&I=K,\end{array}\right$.
for $I,K\in G_{l}^{0}$. A multiple of this matrix is used in the construction of the classical CTQWs on graphs, see, e.g. [5], [4].
12.2. Discretizations of differential equations
We now discuss the discretization of (9.5). Since $u\left(\cdot,t\right)\in\mathcal{C}(\mathbb{Z}_{p})$ for $t\geq 0$, and $\mathcal{D}\left(\mathbb{Z}_{p}\right)$ is dense in $\mathcal{C}(\mathbb{Z}_{p})$, then for $l$ sufficiently large, the function
$u^{\left(l\right)}\left(x,t\right)={\displaystyle\sum\limits_{I\in G_{l}}}u_{I}^{\left(l\right)}\left(t\right)\Omega\left(p^{l}\left|x-I\right|_{2}\right)\text{, with }u_{I}^{\left(l\right)}\left(t\right)\in\mathcal{C}^{1}\left[0,\infty\right)$
is a good approximation in the norm $\left\|\cdot\right\|_{\infty}$ for each fixed $t\geq 0$. By identifying the function $u^{\left(l\right)}\left(x,t\right)$ with the column vector $\left[u_{I}^{\left(l\right)}\right]$, we have the following discretization of (9.5):
(12.5) $\left\{\begin{array}[c]{l}\left[u_{I}^{\left(l\right)}\left(t\right)\right]\in\mathbb{R}^{p^{l}}\text{, }t\geq 0;\text{ }u_{I}^{\left(l\right)}\left(t\right)\in C^{1}(\mathbb{R}_{+})\text{, }I\in G_{l}\\ \\ \frac{\partial}{\partial t}\left[u_{I}^{\left(l\right)}\left(t\right)\right]=\boldsymbol{J}^{\left(l\right)}\left[u_{I}^{\left(l\right)}\left(t\right)\right]\text{, }I\in G_{l},t\geq 0\\ \\ \left[u_{I}^{\left(l\right)}\left(0\right)\right]=\left[u_{I}^{\left(\text{init}\right)}\right]\in\mathbb{R}^{p^{l}}.\end{array}\right$.
In similar way, the discretization of (11.3) is given by
(12.6) $\left\{\begin{array}[c]{l}\left[\Psi_{I}^{\left(l\right)}\left(t\right)\right]\in\mathbb{C}^{p^{l}}\text{, }t\geq 0;\text{ }\Psi_{I}^{\left(l\right)}\left(t\right)\in C^{1}(\mathbb{R}_{+})\text{, }I\in G_{l}\\ \\ \mathrm{i}\frac{\partial}{\partial t}\left[\Psi_{I}^{\left(l\right)}\left(t\right)\right]=-\boldsymbol{J}^{\left(l\right)}\left[\Psi_{I}^{\left(l\right)}\left(t\right)\right]\text{, }I\in G_{l},t\geq 0\\ \\ \left[\Psi_{I}^{\left(l\right)}\left(0\right)\right]=\left[\Psi_{I}^{\left(\text{init}\right)}\right]\in\mathbb{C}^{p^{l}},\end{array}\right$.
see [26], and the references therein, for a further discussion.
The construction of CTQWs on graphs introduced in [2], see, also [5], [4] is based on the equations (12.5)-(12.6), when $\boldsymbol{J}^{\left(l\right)}$ is as in (12.4).
12.3. Quantum networks
From now on, $\mathcal{H}_{l}$ denotes the Hilbert space $\mathbb{C}^{p^{l}}$, with inner product $\left\langle\cdot,\cdot\right\rangle_{l}$, and canonical basis as $\left\{e_{I}\right\}_{I\in G_{l}}$. Since $\boldsymbol{J}^{\left(l\right)}$ is a Hermitian matrix so $\exp(it\boldsymbol{J}^{\left(l\right)})$ is unitary matrix. We define the transition probability $\pi_{I,J}\left(t\right)$ from $J$ to $I$ as
(12.7) $\pi_{I,J}\left(t\right)=\left|\left\langle e_{I}\right|\mathrm{e}^{\mathrm{i}t\boldsymbol{J}^{\left(l\right)}}\left|e_{J}\right\rangle_{l}\right|^{2}\text{, for }J,I\in G_{l}.$
Then
${\displaystyle\sum\limits_{I\in G_{l}}}\pi_{I,J}\left(t\right)={\displaystyle\sum\limits_{I\in G_{l}}}\left|\left\langle e_{I}\right|\mathrm{e}^{\mathrm{i}t\boldsymbol{J}^{\left(l\right)}}\left|e_{J}\right\rangle\right|^{2}=1,$
see [26], for a further details. We call to the continuous-time Markov chain on $G_{l}$ determined by the transition probabilities $\left[\pi_{I,J}\left(t\right)\right]_{I,J\in G_{l}}$, the quantum network associated with the discrete $p$-adic Schrödinger equation (11.3). When the matrix $\boldsymbol{J}^{\left(l\right)}$ has the form (12.4), we recover the CTQWs introduced in [2], see, also [5], [4].
12.4. Additional remarks
We now show that the CTQMC defined in (12.7) agrees with the one given in Theorem 11.2. First, we notice that $\left(\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right),\left\langle\cdot,\cdot\right\rangle\right)$ is isometric to $\left(\mathcal{H}_{l},\left\langle\cdot,\cdot\right\rangle_{l}\right)$. The isometry is defined by
$p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)\rightarrow e_{r}\text{, for }r\in G_{l}.$
Using the results give in Remark 12.2, and the fact that $L^{2}(\mathbb{Z}_{p})\subset L^{1}(\mathbb{Z}_{p})$, there is a projection
$\boldsymbol{P}_{l}:L^{2}\left(\mathbb{Z}_{p}\right)\rightarrow\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right),$
satisfying $\boldsymbol{P}_{l}=\boldsymbol{P}_{l}^{2}$. For $f\in L^{2}\left(\mathbb{Z}_{p}\right)$, $\varphi_{l}=\boldsymbol{P}_{l}f$ is a good approximation of $f$ in $\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)$. The CTQMCs constructed in Theorem 11.2 use the semigroup
$\boldsymbol{P}_{l}\mathrm{e}^{\mathrm{i}t\boldsymbol{J}}:\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)\rightarrow\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right),$
while the ones introduced in this section use the semigroup
$\mathrm{e}^{\mathrm{i}t\boldsymbol{J}^{\left(l\right)}}:\mathcal{H}_{l}\rightarrow\mathcal{H}_{l}.$
After identifying $\mathcal{H}_{l}$ and $\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)$, and using the fact that $\boldsymbol{J}^{\left(l\right)}=\boldsymbol{P}_{l}\boldsymbol{J}$ $\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)\rightarrow\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)$, we have
$\boldsymbol{P}_{l}\mathrm{e}^{\mathrm{i}t\boldsymbol{J}}=\mathrm{e}^{\mathrm{i}t\boldsymbol{P}_{l}\boldsymbol{J}}=\mathrm{e}^{\mathrm{i}t\boldsymbol{J}^{\left(l\right)}}.$
This last equality follows from $\boldsymbol{P}_{l}\boldsymbol{J}^{k}=\left(\boldsymbol{P}_{l}\boldsymbol{J}\right)^{k}$, $k\in\mathbb{N}$, which in turn is a consequence of $\boldsymbol{P}_{l}=\boldsymbol{P}_{l}^{2}$, by using that
$\mathrm{e}^{\mathrm{i}t\boldsymbol{P}_{l}\boldsymbol{J}}={\displaystyle\sum\limits_{k=0}^{\infty}}\frac{\left(it\right)^{k}}{k!}\left(\boldsymbol{P}_{l}\boldsymbol{J}\right)^{k}={\displaystyle\sum\limits_{k=0}^{\infty}}\frac{\left(it\right)^{k}}{k!}\boldsymbol{P}_{l}\boldsymbol{J}^{k}=\boldsymbol{P}_{l}{\displaystyle\sum\limits_{k=0}^{\infty}}\frac{\left(it\right)^{k}}{k!}\boldsymbol{J}^{k}=\boldsymbol{P}_{l}\mathrm{e}^{\mathrm{i}t\boldsymbol{J}}.$
In the last calculation, we used the fact that all the operators involved are bounded.
Remark 12.4.
The above reasoning also shows that in the case
$\boldsymbol{P}_{l}:L^{1}\left(\mathbb{Z}_{p}\right)\rightarrow\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right),$
with $\boldsymbol{P}_{l}=\boldsymbol{P}_{l}^{2}$, we have
$\boldsymbol{P}_{l}\mathrm{e}^{t\boldsymbol{J}}=\mathrm{e}^{t\boldsymbol{P}_{l}\boldsymbol{J}}=\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}:\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)\rightarrow\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right).$
This implies that in the case $\mathbb{J}=G_{l}$, and $\mathcal{K}_{r}=B_{-l}(r)=r+p^{l}\mathbb{Z}_{p}$,
$p_{r,v}(t)$ $=p^{l}{\displaystyle\int\limits_{v+p^{l}\mathbb{Z}_{p}}}Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)dx=p^{l}{\displaystyle\int\limits_{v+p^{l}\mathbb{Z}_{p}}}\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)dx$
$={\displaystyle\int\limits_{\mathbb{Z}_{p}}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)dx,$
for $r,v\in G_{l}$.
13. Discussion
By $p$-adic QM, we mean quantum mechanics, in the sense of Dirac-von Neumann, with quantum states from a complex Hilbert space of the form $L^{2}(B)$, where $B$ is an open compact subset of $\mathbb{Q}_{p}^{N}$. In this framework, the time is a real variable, but the positions are vectors from $\mathbb{Q}_{p}^{N}$. The field of $p$-adic number $\mathbb{Q}_{p}$ is a paramount prototype of a “discrete space:” any continuous trajectory from $\mathbb{R}$ in to $\mathbb{Q}_{p}$ is constant. The motion in $\mathbb{Q}_{p}$ consists of a sequence of jumps. This assertion is also valid in any subset $B\subset\mathbb{Q}_{p}^{N}$; consequently, there are world lines in $B$, and so $p$-adic QM is not compatible with the theory of relativity.
$p$-Adic QM is a model of the Dirac-von Neumann formalism on a discrete space. Here, it is relevant to mention that there are several different types of $p$-adic QM. In some models, time is a $p$-adic variable; this type of model is outside the scope of the Dirac-von Neumann formalism. $p$-Adic QM with $p$-adic time was introduced in the 1980s by Vladimirov and Volovich [18]; the literature on it is extensive. Since the 1980s, a central problem has been to determine whether $p$-adic QM has a physical meaning.
In [39], the author and Mayes studied a $p$-adic version of the infinite potential well. This model describes the confinement of a particle in a $p$-adic ball. They rigorously solved the Cauchy problem for the Schrödinger equation and determined the stationary solutions. By dividing a $p$-adic ball into a finite number of sub-balls and using the wavefunctions of the infinite potential well, they constructed a continuous-time quantum walk (CTQW) on a fully connected graph, where each vertex corresponds to a sub-ball in the partition of the original ball. This approach shows that one can associate a CTQW with any $p$-adic Schrödinger equation. But this approach requires explicit formulas for solving the Cauchy problem for the Schrödinger equation, which is a very involved task and probably impossible for general equations, and, on top of that, the CTQWS produced do not include the ones used in quantum computing, [2]-[5]. Here, we recall that CTQWS are a particular case of continuous-time quantum Markov chains (CTQMCs). In the literature, CTQWs are usually constructed using adjacency matrices, whereas CTQMCs do not require them.
In [26], the author showed that a large class of $2$-adic Schrödinger equations is the scaling limit of CTQMCs. As a practical result, we construct new types CTQWs on graphs using two symmetric matrices. This construction includes, as a particular case, the CTQWs constructed using adjacency matrices. This paper did not include simulations of the CTQMCs, which are essential for understanding their dynamics.
The goal of this paper is to provide a draft of the theory of CTMCs associated with the $p$-adic Schrödinger equations, which contains as a particular case of the CTQWs constructed using adjacency matrices. Here, we study the CTMCs associated with certain $p$-adic heat equations, and the CTQMCs related to the Schrödinger equations obtained by Wick rotation. We solve the initial-value problems associated with the mentioned equations and provide numerical simulations for the corresponding CTQMCs. Behind the results presented in this paper is the conjecture, already formulated in [26], that general $p$-adic heat equations and the corresponding Schrödinger equations are scaling limits of CTMCs, respectively CTQMCs.
The connection between $p$-adic Schrödinger equations and CTQMCs demonstrates that $p$-adic QM has physical meaning. The $p$-adic Schrödinger operators are non-local, so from the beginning, ”spooky actions at a distance” are allowed. Then, interpreting the experimental confirmation of the violation of Bell’s inequalities as ”the universe is non-locally real,” which means that the need to choose between realism and non-locality implies that $p$-adic QM admits realism. The implications of this assertion are explored by the author in [43]; see also [24], [25] for the discussion of other $p$-adic models in QM.
Finally, it is worth noting that the developed framework remains valid if we replace the field of $p$-adic numbers with a more general ultrametric field, while keeping the state space as a Hilbert space of complex-valued functions. If the wavefunctions are not complex-valued, we cannot use the Dirac-von Neumann formalism. In particular, the study of CTQWs with $p$-adic transition probabilities is out of the scope of this work.